### v1

# Using nonassociative algebras to classify skew polycyclic codes up to isometry and equivalence

Abstract.
We propose new definitions of equivalence and isometry for skew polycyclic codes that will lead to tighter classifications than existing ones. This helps to reduce the number of previously known isometry and equivalence classes, and state precisely when these different notions coincide. In the process, we classify classes of skew $(f,\sigma,\delta)$-polycyclic codes with the same performance paramenters, to avoid duplicating already existing codes.
We exploit that the generator of a skew polycyclic code is in one-one correspondence with the generator of a principal left ideal in its ambient algebra. Algebra isomorphisms that preserve the Hamming distance (called isometries) map generators of principal left ideals to generators of principal left ideals and preserve length, dimension and Hamming distance of the codes. We allow the ambient algebras to be nonassociative, thus eliminating the need on restrictions on the length of the codes.
The isometries between the ambient algebras can also be used to classify corresponding linear codes equipped with the rank metric.
Key words and phrases:
skew cyclic codes, skew polycyclic codes, quantum error-correcting codes, nonassociative algebra, isometry
2020 Mathematics Subject Classification:
Primary: 94B40, 94B05; Secondary: 17A35, 94B05
Introduction
Any results on equivalences for skew polycyclic codes can be used to optimize searches for good codes, avoid duplicating existing code parameters, and find equivalence classes in given sets of codes. Therefore there has been extensive work towards classifying constacyclic, polycyclic, skew constacyclic and more generally skew polycyclic codes (also called skew $(\sigma,\delta)$-polycyclic codes or $(\sigma,\delta)$-codes in [BouUl2009, BouLe2013, BouUl2014]) via different equivalence relations, exploiting their algebraic properties, see for example [CFLL2012, CDL2015, CON2025, OaNA25]. For instance, in [BLMS2024] left submodules of $R/Rf$ represent the skew skew $(f,\sigma,\delta)$-polycyclic codes and $(\sigma,\delta)$-pseudo-linear transformations when $f$ is not two-sided, and the Hamming isometrical equivalent $(\sigma,\delta)$-polycyclic codes are classified.
Let $S$ be a unital commutative associative ring and $f\in R=S[t;\sigma,\delta]$ a monic skew polynomial. When the context is clear, we will simply write skew polycyclic code instead of skew $(f,\sigma,\delta)$-polycyclic code, and skew constacyclic code instead of skew $(\sigma,c)$-constacyclic code (which is the when $f(t)=t^{m}-a$).
In the present paper, we extend and unify the existing classifications employing isometries between nonassociative algebras, as initiated in [NevPu2025].
1. We view skew $(f,\sigma,\delta)$-polycyclic codes as left principal ideals in their ambient algebras $R/Rf=S[t;\sigma,\delta]/S[t;\sigma,\delta]f$ [Pum2017]. This approach is well known when $f$ generates a two-sided ideal in $R$, e.g. when $f$ lies in the center of $R$. For instance, some results for skew constacyclic codes over fields developed in [LM25] were given using the associative ambient algebras $K[t;\sigma]/K[t;\sigma](t^{m}-a)$. When $Rf$ is not a two-sided ideal, the $R$-module $R/Rf$ still carries a ring structure, but a nonassociative one, and is a nonassociative Petit algebra over subrings of its center.
A large percentage of the existing results on skew polycyclic codes, when viewed as ideals in associative algebras, rely on multiplications of not more than two elements in the ideal, or can be effortlessly adjusted to the nonassociatie setting. This basically guarantees their validity also when the ambient algebra is not associative.
2. We call two skew polycyclic codes isometric if there is a Hamming-weight preserving map between them that is canonically induced by a monomial isomorphism of degree $k$ between their ambient Petit algebras mapping $t$ to $\alpha t^{k}$ for some invertible $\alpha\in S^{\times}$ and restricting to some $\tau\in{\rm Aut}(S)$ over $S$, viewed as algebras over some subring of their center, where the Hamming weight of a polynomial $p(t)$ is defined as the number of nonzero coefficients of $p(t)$. Any such an isomorphism between Petit algebras, called an isometry, gives a correspondence between their ideals, hence between the associated codes, which preserves code rate and Hamming distance as well as length and dimension. This approach generalizes the previously used notions of isometry and equivalence between skew polycyclic codes that were defined using algebraic structures so far (which exclusively worked with isometries restricting to some $id\in{\rm Aut}(S)$), and is lead by our knowledge of isomorphisms between nonassociative Petit algebras [BrPum2019, Br2018, BrPumSt2018, BrPum2018, Pum2024].
3. We compute all the classes of codes that are equivalent to a given one (Theorem 5.6). Our refined notion of equivalence yields tighter classifications, for some first examples for codes over finite fields see Section 5.2. (This was observed for the first time in [NevPum2025, Example 5.8].) We give a list of new and easy necessary and sufficient criteria for isometry and equivalence of skew polycyclic codes that will help distinguish non-equivalent codes with potentially different performance parameters.
We will focus mostly on $(f,\sigma)$-polycyclic codes, especially skew constacyclic codes, as for these the isomorphisms and automorphisms of the corresponding ambient Petit algebras are well-known. By allowing the case that $\sigma=id$, we include the case that $R=S[t]$ where possible, thus recovering results from [CFLL2012, CDL2015, OuaHA25] as special cases. Overall, our equivalence relations will result in tighter partitions of non-equivalent classes of codes than the ones defined in [ALY2022, CON2025, OaNA25].
(Skew) polycyclic codes over finite rings, e.g. over mixed alphabets, are among the most prominent linear codes currently studied. They can be used to construct quantum error-correcting codes, usually via Gray maps and the CSS construction, see for instance [BagPa2025, YSP23, BDUBY, VPSI23, PPI25]. We expect that the results of this paper will make the search for new SPC codes and correspondingly quantum error-correcting codes more efficient and more organized. Good parameters for quantum codes are still under investigation, and any results on isometries and equivalences for SPC codes can be used to optimize search algorithms, in turn facilitating the discovery of new record-breaking quantum codes. In particular, when considering the CSS construction for quantum error-correcting codes, one needs to concentrate on codes $C$ such that the dual code $C^{\perp}$ satisfies $C^{\perp}\subset C$ as explained in [DinhBagUpadhyayBandiTansuchat2021]. Thus an important next step in the classification will be to consider when isometries preserve some suitable notion of duality of codes.
Our main results can be found in Section 5 onwards and employ Hamming weight preserving isomorphisms between Petit algebras treated in Section 4. We look at the case that $\delta=0$ and $\sigma$ an automorphism.
Given two monic polynomials
$f(t)=t^{m}-\sum_{i=0}^{m-1}a_{i}t^{i},\quad h(t)=t^{m}-\sum_{i=0}^{m-1}b_{i}t^{i}\in R=S[t;\sigma],$
we call two classes of skew $(f,\sigma)$ and $(h,\sigma)$-polycyclic codes of length $m$ isometric if there is an algebra isomorphism $G_{\tau,\alpha,k}$ between their ambient algebras $R/Rf$ and $R/Rh$ which restricts to some $\tau\in{\rm Aut}(S)$ and maps $t$ to $\alpha t^{k}$ for some integer $k$ with 0<k<m10<k<m-1 and some $\alpha\in S^{\times}$, and Chen isometric if $\tau=id$. When $k=1$, we call the classes equivalent, and Chen equivalent, when additionally $\tau=id$, and write $G_{\tau,\alpha}=G_{\tau,\alpha,k}$. The current literature only uses Chen isometry and Chen equivalence.
Theorem.
(Theorem 5.1) Let $\alpha\in S^{\times}$ and assume that $\tau\in{\rm Aut}(S)$ commutes with $\sigma$. Then the two classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes of length $m$ over $S$ are equivalent if and only if there exists $\tau\in{\rm Aut}(S)$ that commutes with $\sigma$ and $\alpha\in S^{\times}$ such that
$\tau(a_{i})=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$
for all $i\in\{0,\ldots,m-1\}$ (resp., Chen-equivalent if and only if above equation is true for $\tau=id$).
For codes over fields, isometry and equivalence coincide in a large number of cases.
Theorem.
(Theorem 5.2) Suppose $K$ is a field and $\mathrm{Aut}(K)$ is abelian. Let $n\geq m-1$ and assume that $f$ and $h$ do not generate two-sided ideals in $K[t;\sigma]$. Then two classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes of length $m$ over $S$ are equivalent if and only if they are isometric, and Chen equivalent if and only if they are Chen-isometric.
Note that it is known that if $S_{0}$ denotes the ring fixed by $\sigma$, and $a,b\in S\setminus S_{0}$, or $a,b\in S_{0}^{\times}$ and $n\nmid m$ then equivalence and isometry coincide for the classes of skew $(\sigma,a)$-constacyclic and skew $(\sigma,b)$-constacyclic codes of length $m$ [NevPu2025, Corollary 4.5].
The equivalence classes ${\bf C}_{h}]$ of classes of skew polycyclic codes of length $m$ over $S$ that are equivalent to ${\bf C}_{h}$ (respectively, Chen equivalent) are computed in Theorem 5.6. For finite rings we count the number of Chen isometry classes of skew constacyclic codes for classes of codes with a nonassociative ambient algebra (Theorem 4.11).
Now there is an intricate connection between classes of equivalent skew $(\sigma,f)$-polycyclic and equivalent skew $(\sigma,a)$-constacyclic codes over commutative rings $S$.
Two classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes over $S$ of length $m$ are equivalent via $G_{\tau,\alpha}$ if and only if for all $i\in\{0,\ldots,m-1\}$ where $a_{i}\not=0$, the class of $(\sigma,a_{i})$-constacyclic codes is (mi)(m-i)-equivalent to the class of $(\sigma,b_{i})$-constacyclic codes via $G_{\tau,\sigma^{i}(\alpha)}$ (Proposition 5.7).
For $a,b\in S^{\times}$, the classes of skew $(\sigma,a)$-constacyclic and skew $\sigma$-cyclic codes of length $m$ are equivalent if and only if $a\in N_{m}^{\sigma}(S^{\times}).$ The classes of skew $(\sigma,a)$-constacyclic codes and skew $\sigma$-negacyclic codes of length $m$ are equivalent if and only if $a\in N_{m}^{\sigma}(S^{\times}).$ In both cases, isometry and equivalence coincide when $m$ does not divide $n$ (Corollary 5.10).
Conditions for when two skew polycyclic codes cannot be equivalent, are given in Section 5.3.
Our results for isometries between the ambient algebras also have consequences for classifying skew polycyclic codes with respect to the rank metric, especially over finite fields $K$ (where we need $\alpha\in K^{\prime}$, the suitable subfield of $K$ we consider the rank metric over). In particular, they will help understand the different MRD codes obtained from these [OaHA25, Proposition 3.13]. The most canonical approach here will be to additionally employ the isometries $G_{\tau,\alpha}$, for all $\tau\in{\rm Aut}(K)$ (Remark 3.6).
The readers interested in codes are encouraged to skip the technical sections dealing with isomorphisms between nonassociative algebras, and focus on the results achieved for skew polycyclic codes.
1. Preliminaries
1.1. Nonassociative rings
Let $R$ be a unital commutative ring and let $A$ be an $R$-module. We call $A$ a (nonassociative) algebra over $R$ if there exists an $R$-bilinear map $A\times A\mapsto A$, $(x,y)\mapsto x\cdot y$, usually denoted simply by juxtaposition $xy$, the multiplication of $A$. An algebra $A$ is called unital if there is an element in $A$, denoted by 1, such that $1x=x1=x$ for all $x\in A$. We will only consider unital algebras.
If we “forget” the $R$-module structure of $A$ then $A$ is a nonassociative ring, with $(A,+)$ being an abelian additive group and $(A,\cdot)$ a groupoid.
Define [x,y,z]=(xy)zx(yz)[x,y,z]=(xy)z-x(yz). The associativity in $A$ is measured by the left nucleus defined as ${\rm Nuc}_{l}(A)=\{x\in A\,|\,[x,A,A]=0\}$, the middle nucleus ${\rm Nuc}_{m}(A)=\{x\in A\,|\,[A,x,A]=0\}$ and the right nucleus ${\rm Nuc}_{r}(A)=\{x\in A\,|\,[A,A,x]=0\}$. The sets ${\rm Nuc}_{l}(A)$, ${\rm Nuc}_{m}(A)$ and ${\rm Nuc}_{r}(A)$ are associative subalgebras of $A$. Their intersection ${\rm Nuc}(A)=\{x\in A\,|\,[x,A,A]=[A,x,A]=[A,A,x]=0\}$ is the nucleus of $A$ and is an associative subalgebra of $A$ containing $R1$ and $x(yz)=(xy)z$ whenever one of the elements $x,y,z$ is in ${\rm Nuc}(A)$. The commuter of $A$ is defined as ${\rm Comm}(A)=\{x\in A\,|\,xy=yx\text{ for all }y\in A\}$ and the center of $A$ is ${\rm C}(A)=\text{Nuc}(A)\cap{\rm Comm}(A)$ [Sch1995].
A nonassociative algebra $A\not=0$ (resp., a ring $A\not=0$) over a field $F$ is called a division algebra, resp., a division ring, if for all $a\in A$, $a\not=0$, the left and right multiplication with $a$, $L_{a}(x)=ax$ and $R_{a}(x)=xa$, are bijective maps. If $A$ is a finite-dimensional algebra over $F$, then $A$ is a division algebra if and only if $A$ has no zero divisors [Sch1995].
A nonassociative ring or algebra $A$ is called a proper nonassociative ring or algebra if it is not associative.
Two $F$-algebras AA and $A^{\prime}$ are called $F$-isotopic if and only if there exist $F$-linear isomorphisms $F,G,H:A\longrightarrow A^{\prime}$ such that $H(uv)=F(u)G(v)$ for all $u,v\in A$.
1.2. Skew polynomial rings
Let $S$ be a unital associative ring (it need not be commutative), and $\sigma$ a ring automorphism of $S$. A left $\sigma$-derivation is an additive map $\delta:S\rightarrow S$, such that $\delta(ab)=\sigma(a)\delta(b)+\delta(a)b$ for all $a,b\in S$. Define $S_{0}={\rm Fix}(\sigma)=\{a\in S\,|\,\sigma(a)=a\}$ and ${\rm Const}(\delta)=\{a\in S\,|\,\delta(a)=0\}$.
The skew polynomial ring $R=S[t;\sigma,\delta]=\left\{a_{0}+a_{1}t+\dots+a_{n}t^{n}\,|\,a_{i}\in S,n\in\mathbb{N}\right\}$ is the set of skew polynomials $a_{0}+a_{1}t+\dots+a_{n}t^{n}$, $a_{i}\in S$, with the usual addition of two polynomials, where the multiplication is defined by the rule $ta=\sigma(a)t+\delta(a)$ for all $a\in S$. That means
$at^{n}bt^{m}=\sum_{j=0}^{n}a(\Delta_{n,j}\,b)t^{m+j}$
for all $a,b\in S$, where the map $\Delta_{n,j}$ is defined recursively via
$\Delta_{n,j}=\delta(\Delta_{n-1,j})+\sigma(\Delta_{n-1,j-1}),$
with $\Delta_{0,0}=id_{S}$, $\Delta_{1,0}=\delta$, $\Delta_{1,1}=\sigma$ [Ore1933]. Therefore $\Delta_{n,j}$ is the sum of all monomials in $\sigma$ and $\delta$ of degree $j$ in $\sigma$ and degree njn-j in $\delta$ [Jac1996, p. 2]. If $\delta=0$, then $\Delta_{n,n}=\sigma^{n}$.
When $\sigma=id$ and $\delta=0$, $S[t]=S[t;id,0]$ is the usual ring of left polynomials. When $\delta=0$, we use the notation $S[t;\sigma]=S[t;id,0]$.
For $f(t)=a_{0}+a_{1}t+\dots+a_{n}t^{n}\in R$ with $a_{n}\not=0$ define ${\rm deg}(f)=n$ and ${\rm deg}(0)=$-$\infty$. Then ${\rm deg}(gh)\leq{\rm deg}(g)+{\rm deg}(h)$ (with equality if $h$ or $g$ have an invertible leading coefficient, or if $S$ is a division ring). An element $f\in R$ is irreducible in $R$ if it is not a unit and it has no proper factors, i.e if there do not exist $g,h\in R$ with $1\leq{\rm deg}(g),{\rm deg}(h)<{\rm deg}(f)$ such that $f=gh$.
Two nonzero skew polynomials $f$ and $g$ in $R$ are similar, written $f\sim g$, if $R/Rf\cong R/Rg$ as left $R$-modules, equivalently, there exist $u,v\in R$, such that $gcrd(f,u)=1$, $gcld(g,v)=1$ and $gu=vf$ [Jac1996, p. 11]. The element $u\in R$ can be chosen such that $u\in R_{m}$ [Jac1996, p. 15].
We call $f\in R$ two-sided if $Rf$ is a two-sided ideal in $R$.
The companion matrix of a monic polynomial $f=t^{m}-\sum_{i=0}^{m-1}a_{i}t^{i}\in R=S[t;\sigma,\delta]$ of degree $m$ is given by
$A_{f}=\left(\begin{array}[]{cccccccc}0&1&\cdots&&&&0\\ &0&1\cdots&&&\cdots&\\ \vdots&&\ddots&&&&\vdots\\ &&&\ddots&&&\\ &&&&\ddots&&\\ \vdots&&&&&\ddots&1\\ a_{0}&a_{1}&\cdots&&&\cdots&a_{m-1}\\ \end{array}\right)$
Let $h\in R=S[t;\sigma,\delta]$ be a monic polynomial then $f\sim h$ if and only if there exists an invertible matrix $B\in M_{m}(K)$ such that $A_{f}=\sigma(B)A_{h}B^{-1}+\delta(B)$. Here, $\sigma$ respectively $\delta$ applied to $B$ means applying it to each entry of the matrix.
Note that the fact that $R/Rf\cong R/Rh$ as left $R$-modules does not imply that $R/Rf\cong R/Rh$ as nonassociative rings or algebras. In certain cases, however, it may imply that $R/Rf$ and $R/Rh$ are isotopic rings [NevPum2025, Theorem 3.4].
1.3. Isomorphisms between skew polynomial rings
Let $S$ be a unital associative ring (it need not be commutative), $f,h\in R=S[t;\sigma,\delta]$ monic. For all $\tau\in{\rm Aut}(S)$ and $\alpha,\beta\in S$ with $\alpha$ invertible, such that
$\beta\tau(b)+\alpha\delta(\tau(b))=\tau(\sigma(b))\beta+\tau(\delta(b))\text{ and }\alpha\sigma(\tau(b))=\tau(\sigma(b))\alpha$
for all $b\in S$, the map $H_{\tau,\beta,\alpha}:S[t;\sigma,\delta]\longrightarrow S[t;\sigma,\delta]$,
$H_{\tau,\beta,\alpha}:S[t;\sigma,\delta]\longrightarrow S[t;\sigma,\delta],\quad\sum_{i=0}^{n}b_{i}t^{i}\mapsto\sum_{i=0}^{n}\tau(b_{i})(\beta+\alpha t)^{i}$
is a ring automorphism of $R$ [LaLe92, p. 4]. In particular, if
$\alpha\delta(\tau(b))=\tau(\delta(b))\text{ and }\alpha\sigma(\tau(b))=\tau(\sigma(b))\alpha$
for all $b\in S$ then
$G_{\tau,\alpha}:S[t;\sigma,\delta]\longrightarrow S[t;\sigma,\delta],$
$\sum_{i=0}^{n}b_{i}t^{i}\mapsto\sum_{i=0}^{n}\tau(b_{i})(\alpha t)^{i}$
is a ring automorphism that preserves the Hamming weight.
We note that the ring isomorphisms $G=G_{\tau,\alpha}:R/Rf\rightarrow R/Rh$ for some $\alpha\in S^{\times}$, are $S^{\prime}$-algebra isomorphisms, where $S^{\prime}\subset{\rm Fix}(\tau)\cap{\rm Const}(\delta)\cap{\rm Fix}(\sigma)\cap C(S)$.
For the special case that $\delta=0$ we know by [Rim1978, Theorem 3] that the map $G:S[t;\sigma]\rightarrow S[t;\sigma]$, $G(t)=\sum_{i=0}^{m-1}\alpha_{i}t^{i}$, $\alpha_{i}\in S$, extends $\tau\in{\rm Aut}(S)$ to an isomorphism $G:S[t;\sigma]\rightarrow S[t;\sigma]$ if and only if
  1. (1)
    $\sigma(\tau(a))k_{i}=k_{i}\tau(\sigma^{i}(a))$ for all $a\in S$ and $i\in\{0,\dots,m\},$
  2. (2)
    $\alpha_{1}\in S^{\times},$
  3. (3)
    $\alpha_{i}\text{ is nilpotent for all }i\in\{2,\dots m\}.$
Note that if $S$ is a division algebra, for example a field, $\alpha_{i}=0$ for all $i\in\{2,\dots,m\}$ holds automatically.
Since we are only interested in maps that preserve the Hamming distance, we will only consider the case that $\alpha_{0}=0$ and $\alpha_{i}=0$ for all $i\in\{2,\dots,m\}$. We conclude that $G_{\tau,\alpha}:S[t;\sigma]\rightarrow S[t;\sigma]$ extends $\tau\in{\rm Aut}(S)$ to an isomorphism that respects the Hamming weight with $G(t)=\alpha t$ for some $\alpha\in K^{\times}$, if and only if $\sigma(\tau(a))\alpha=\alpha\tau(\sigma(a))$ for all $a\in S$, and $\alpha\in S^{\times}$
2. Skew polycyclic codes viewed as ideals in Petit algebras
2.1. Petit algebras
Let $f\in R=S[t;\sigma,\delta]$ have an invertible leading coefficient. Then for all $g\in R$ of degree $l\geq m$, there exist uniquely determined $r,q\in R$ with ${\rm deg}(r)<{\rm deg}(f)$, such that $g(t)=q(t)f(t)+r(t),$ and since we assume $\sigma\in{\rm Aut}(S)$, also uniquely determined $r,q\in R$ with ${\rm deg}(r)<{\rm deg}(f)$, such that $g(t)=f(t)q(t)+r(t)$ [Pum2017, Proposition 1].
Let ${\rm mod}_{r}f$ denote the remainder of right division by $f$ and ${\rm mod}_{l}f$ the remainder of left division by $f$. Since the remainders are uniquely determined, the set $\{g\in R\,|\,{\rm deg}(g)<m\}$ canonically represent the elements of the (left, resp. right) $R$-modules $R/Rf$ and $R/fR$.
Theorem 2.1.
$(i)$ The additive group $\{g\in R\,|\,{\rm deg}(g)<m\}$ together with the multiplication
$gh=gh\,\,{\rm mod}_{r}f,$
where the right hand side is the remainder of $gh$ after division by $f$ on the right, is a unital nonassociative ring denoted by denoted by $R/Rf$ or $S_{f}$.
$(ii)$ The additive group $\{g\in R\,|\,{\rm deg}(g)<m\}$ together with the multiplication
$gh=gh\,\,{\rm mod}_{l}f,$
where the right hand side is the remainder of $gh$ after division by $f$ on the left, is a unital nonassociative ring denoted by $R/fR$ or $_{f}S$.
Note that $S_{f}$ is associative exactly when $S[t;\sigma,\delta]f$ is a two-sided ideal. For example, this happens when $S=K$ is a field, $f(t)=t^{m}-\sum_{i=0}^{m-1}a_{i}t^{i}\in S_{0}[t]$ and $\sigma^{m-i}=id$ for all those $i$ where $a_{i}\not=0$ [Pe1966, (15)].
Without loss of generality we will only consider monic polynomials, since $S_{f}=S_{af}$ for all invertible $a\in S$. If $f$ is reducible then $S_{f}$ contains zero divisors. In the following, we will thus always assume that
[…]
Both $S_{f}$ and $_{f}S$ are unital nonassociative algebras (called Petit algebras) over the commutative subring
$\{a\in S\,|\,ah=ha\text{ for all }h\in S_{f}\}={\rm Comm}(S_{f})\cap S$
of $S$, and have center
$C(S)\cap{\rm Fix}(\sigma)\cap{\rm Const}(\delta).$
If $S$ is a division ring and $S_{f}$ is not associative then $C(S_{f})=\{a\in S\,|\,ah=ha\text{ for all }h\in S_{f}\}$.
In the following, we call the algebras $S_{f}$ and $_{f}S$ Petit algebras [Pe1966] (the original paper only considered division rings $S$). We will focus on the algebras $S_{f}$, since $_{f}S$ is anti-isomorphic to $S_{f}$. We note that the principal right ideals of the algebras $_{f}S$ play an equally relevant role in linear code constructions along similar arguments as presented in [BLMS2024].
Proposition 2.2.
[Pum2017, Proposition 3] Let $f\in R=S[t;\sigma,\delta]$ have an invertible leading coefficient. The canonical anti-automorphism
$\psi:S[t;\sigma,\delta]\rightarrow S^{op}[t;\sigma^{-1},-\delta\circ\sigma^{-1}],$
$\psi(\sum_{k=0}^{n}a_{k}t^{k})=\sum_{k=0}^{n}(\sum_{i=0}^{k}\Delta_{n,i}(a_{k}))t^{k}$
between the rings $S[t;\sigma,\delta]$ and $S^{op}[t;\sigma^{-1},-\delta\circ\sigma^{-1}]$ induces an anti-automorphism between the nonassociative rings
$S_{f}=S[t;\sigma,\delta]/S[t;\sigma,\delta]f$
and
$_{\psi(f)}S=S^{op}[t;\sigma^{-1},-\delta\circ\sigma^{-1}]/\psi(f)S^{op}[t;\sigma^{-1},-\delta\circ\sigma^{-1}].$
If $\delta=0$ we have
$\psi(\sum_{k=0}^{n}a_{k}t^{k})=\sum_{k=0}^{n}\sigma^{-k}(a_{k})t^{k}.$
For $0\not=a\in S_{f}$, left multiplication $L_{a}$ is an $S_{0}$-module endomorphism. Moreover, right multiplication $R_{a}$ is a left $S$-module homomorphism for all $0\neq a\in S_{f}$.
Let $f\in S[t;\sigma,\delta]$ have degree $m\geq 2$ and an invertible leading coefficient. Then $S_{f}$ is a free left $S$-module of rank $m$ with basis $t^{0}=1,t,\dots,t^{m-1}$. The algebra $S_{f}$ is associative if and only if $Rf$ is a two-sided ideal in $R$. In that case, it coincides with the classical quotient algebra. The $S_{f}$ thus generalizes the classical quotient algebra construction when factoring out a two-sided ideal generated by a two-sided (also called right invariant) skew polynomial $f$.
If $S_{f}$ is not associative then $S\subset{\rm Nuc}_{l}(S_{f}),\,\,S\subset{\rm Nuc}_{m}(S_{f})$ and
$\{g\in R\,|\,{\rm deg}(g)<m\text{ and }fg\in Rf\}={\rm Nuc}_{r}(S_{f}).$
When $S$ is a division ring, these inclusions become equalities.
Proposition 2.3.
[Pum2017] Let $S$ be a unital associative division ring and $f\in R=S[t;\sigma,\delta]$.
$(i)$ All left ideals in $S_{f}$ are generated by some monic right divisor $g$ of $f$ in $R$.
$(ii)$ If $f$ is irreducible, then $S_{f}$ has no non-trivial left ideals.
Example 2.4.
[Pum2017] Let $\sigma\in{\rm Aut}(S)$ have order $n$ and ${\rm Fix}(\sigma)=S_{0}$. For $f(t)=t^{m}-d\in R=S[t;\sigma]$, $d\in S$ invertible, the multiplication in the $S_{0}$-algebra $R/R(t^{m}-d)$ is defined via
$(at^{i})(bt^{j})=\begin{cases}a\sigma^{i}(b)t^{i+j}&\text{if }i+j<m,\\ a\sigma^{i}(b)t^{(i+j)-m}d&\text{if }i+j\geq m,\end{cases}$
for all $a,b\in S$ and then linearly extended. In the literature, $S$ is often a field in which case we know that $R/R(t^{m}-d)$ is an associative algebra over the fixed field of $\sigma$, if and only if $d\in S_{0}$ and $m$ divides the order $n$ of $\sigma$ [Pe1966, (7), (9)]. For $d=1$ in particular, $f$ is two-sided if and only $m|n$.
It is straightforward to see that every ring isomorphism $G:S[t;\sigma,\delta]\rightarrow S[t;\sigma,\delta]$ canonically induces an isomorphism between the algebras $R/Rf$ and $R/RG(f)$, since $G|_{R_{m}}:R_{m}\rightarrow R_{m}$ and $G(g\circ_{f}h)=G(gh-sf)=G(h)G(g)-G(s)G(f)=G(g)\circ_{G(f)}G(h)$ for some $s\in R_{m}$. Note that $G(f(t))=g(t)=\sum_{i=0}^{m}\tau(a_{i})(\alpha t)^{i}\in S[t;\sigma,\delta]$ is usually not monic and can be written as a product of a monic polynomial and some invertible scalar in $S$. We will use this observation repeatedly.
Lemma 2.5.
Let $\alpha\in S$ be invertible, $\tau\in{\rm Aut}(S)$, and $f(t)=t^{m}-\sum_{i=0}^{m-1}a_{i}t^{i}\in S[t;\sigma,\delta]$ a monic polynomial and $h(t)=(\alpha t)^{m}-\sum_{i=0}^{m}\tau(a_{i})(\alpha t)^{i}\in S[t;\sigma,\delta]$. If
$\alpha\delta(\tau(b))=\tau(\delta(b))\text{ and }\alpha\sigma(\tau(b))=\tau(\sigma(b))d$
for all $b\in S$, then there exists an equivalence $G_{\tau,\alpha}:R/Rf\cong R/Rg$.
2.2. Skew $(f,\sigma,\delta)$-polycyclic codes
A left (resp., right) linear code of length $m$ over $S$ is a left (resp., right) submodule of the $S$-module $S^{m}$. From now on, let $f(t)=t^{m}-\sum_{i=0}^{m-1}a_{i}t^{i}\in S[t;\sigma,\delta]$ be a reducible monic polynomial of degree $m>1$.
Definition 2.6.
$(i)$ A left linear code $C\subset S^{m}$ is called a right skew $(f,\sigma,\delta)$-polycyclic code, if for each codeword $(c_{0},c_{1},\ldots,c_{m-1})$ of $C$, also
$\left(0,\sigma(c_{0}),\sigma(c_{1}),\ldots,\sigma(c_{m-2})\right)+\sigma(c_{m-1})(a_{0},a_{1},\ldots,a_{m-1})+(\delta(c_{0}),\delta(c_{1}),\ldots,\delta(c_{m-1}))\in C.$
$(ii)$ A left linear code $C\subset S^{m}$ is called a left skew $(\sigma,\delta,f)$-polycyclic code, if for each codeword $(c_{0},c_{1},\ldots,c_{m-1})$ of $C$, also
$\left(\sigma(c_{1}),\sigma(c_{2}),\ldots,\sigma(c_{m-1}),0\right)+\sigma(c_{0})(a_{0},a_{1},\ldots,a_{m-1})+(\delta(c_{0}),\delta(c_{1}),\ldots,\delta(c_{m-1}))\in C.$
We will only look at right skew polycyclic codes unless explicitly stated otherwise. We can rewrite $(i)$ as
$\left(0,\sigma(c_{0}),\sigma(c_{1}),\ldots,\sigma(c_{m-2})\right)+(\sigma(c_{m-1})a_{0},\sigma(c_{m-1})a_{1},\ldots,\sigma(c_{m-1})a_{m-1})+\delta(c)$
$\left(\sigma(c_{m-1})a_{0},\sigma(c_{0})+\sigma(c_{m-1})a_{1},\sigma(c_{1})+\sigma(c_{m-2})a_{2},\ldots,\sigma(c_{m-2})+\sigma(c_{m-1})a_{m-1})\right)+\delta(c)\in C.$
A skew $(f,\sigma,\delta)$-polycyclic code $C$ over $S$ is called a skew $(\sigma,d)$-constacyclic code, if $\delta=0$ and $f(t)=t^{m}-d$ for a non-zero $d\in S$, that is
$(c_{0},\dots,c_{m-1})\in C\Rightarrow(\sigma(c_{m-1})d,\sigma(c_{0}),\dots,\sigma(c_{m-2}))\in C.$
We will use the bijective map $\Phi:S^{m}\longrightarrow R/Rf$,
$(c_{0},c_{1},\dots,c_{m-1})\mapsto\sum_{i=0}^{m-1}c_{i}t^{i},$
to “move” between linear codes and ideals in the algebras $R/Rf$: For a linear code $C$ of length $m$ over $S$ we denote by $C(t)$ the set of skew-polynomials $h(t)=\sum_{i=0}^{m-1}c_{i}t^{i}$ associated to the codewords $(c_{0},\dots,c_{m-1})\in C$.
Our linear codes $C$ of length $m$ consist of all vectors $(h_{0},\dots,h_{m-1})\in S^{m}$ obtained from the elements $h(t)=\sum_{i=0}^{m-1}h_{i}t^{i}$ in a left principal ideal of $S_{f}$: A skew $(f,\sigma,\delta)$-code $C\subset S^{m}$ is a subset of $S^{m}$ consisting of the vectors $(a_{0},\dots,a_{m-1})$ obtained from elements $h=\sum_{i=0}^{m-1}a_{i}t^{i}$ in a left principal ideal $gS_{f}=S[t;\sigma,\delta]g/S[t;\sigma,\delta]f$ of $S_{f}$, with $g$ a right divisor of $f$ which is assumed to be monic where possible (e.g., when $S$ is a field) and the polynomial of smallest degree that generates the left ideal. This means that the set of vectors corresponding to the elements $\{g,tg,\ldots,t^{k-1}g\}$ in $S_{f}$ forms a basis of $C$ and the dimension of $C$ is $k=n-\deg(g)$. This also means that there is a one-to-one correspondence between principal left ideals of $S_{f}$ and skew $(\sigma,d)$-constacyclic codes of length $n$ over $S$ [Pum2017]. Note that in the setup we study here, $g$ is uniquely determined only up to similarity of skew polynomials.
Theorem 2.7.
[Pum2017] Let $g=\sum_{i=0}^{r}g_{i}t^{i}$ be a monic polynomial which is a right divisor of $f$.
$(i)$ The skew $(f,\sigma,\delta)$-polycyclic code $C\subset S^{m}$ corresponding to the principal ideal $gS_{f}$ is a free left $S$-module of dimension $m-{\rm deg}g$.
$(ii)$ If $(c_{0},\dots,c_{m-1})\in C$ then $L_{t}(c_{0},\dots,c_{m-1})\in C$ (this is the left multiplication with $t$ in the algebra $R/Rf$).
(iii) The matrix generating $C$ represents the right multiplication $R_{g}$ with $g$ in $R/Rf$, calculated with respect to the basis $1,t,\dots,t^{m-1}$, identifying elements $h=\sum_{i=0}^{m-1}h_{i}t^{i}$ with vectors $(h_{0},\dots,h_{m-1})$ via $\Phi$.
Theorem 2.8.
Let $g=\sum_{i=0}^{r}g_{i}t^{i}$ be a monic polynomial which is a right divisor of $f$, such that $f=gh=h^{\prime}g$ for two monic polynomials $h,h^{\prime}\in S[t;\sigma,\delta]$. Let $C$ be the skew $(f,\sigma,\delta)$-polycyclic code corresponding to $g$ and $c=\sum_{i=0}^{m-1}c_{i}t^{i}\in S[t;\sigma,\delta]$. Then the following are equivalent:
$(i)$ $(c_{0},\dots,c_{m-1})\in C$.
$(ii)$ $c(t)h(t)=0$ in $S_{f}$.
(iii) $L_{c}(h)=ch=0$, resp. $R_{h}(c)=hc=0$.
This shows that sometimes $h$ is a parity check polynomial for $C$ also when $f$ is not two-sided. Note that parity check and generator polynomials will only be unique up to similarity in our setting. Also, when we only have $hg=f$, $h$ monic, and $C$ is the code generated by $g$ then if $ch=0$ in $S_{f}$, $c$ is a codeword of $C$.
Corollary 2.9.
Let $g=\sum_{i=0}^{r}g_{i}t^{i}$ be a monic polynomial which is a right divisor of $f$, such that $f=gh=h^{\prime}g$ for two monic polynomials $h,h^{\prime}\in S_{f}$. Let $C$ be the skew $(f,\sigma,\delta)$-polycyclic code corresponding to $g$. Then the matrix representing right multiplication $R_{h}$ with $h$ in $S_{f}$ with respect to the basis $1,t,\dots,t^{m-1}$ is a control matrix of the cyclic $(f,\sigma,\delta)$-code corresponding to $g$.
In particular we obtain:
Theorem 2.10.
Let $f(t)\in S[t;\sigma]$ be monic of degree $m$ and $g=\sum_{i=0}^{r}g_{i}t^{i}\in R/Rf$ be a polynomial which is a right divisor of $f$ (monic if $S$ is a division ring). Let $C\subset S^{m}$ be the skew $(f,\sigma)$-polycyclic code consisting of all $(c_{0},\dots,c_{m-1})$ such that $h(t)=\sum_{i=0}^{m-1}c_{i}t^{i}\in Rg/R$.
$(i)$ If $(c_{0},\dots,c_{m-1})\in C$ then take the left multiplication in $S_{f}$ with $t^{i}$, and conclude that $L_{t^{i}}(c_{0},\dots,c_{n-1})\in C$ for all $1\leq i\leq n-1$.
$(ii)$ Every left ideal generated by $g$ yields a skew $(f,\sigma)$-polycyclic code $C\subset S^{m}$
(iii) The skew $(f,\sigma)$-polycyclic code $C\subset S^{m}$ corresponding to the principal left ideal in $R/Rf$ generated by $g$ is a free left $S$-module of dimension $n-{\rm deg}(g)$.
(iii) The matrix
(2.1) $G=\left(\begin{array}[]{cccccccc}g_{{}_{0}}&g_{{}_{1}}&\cdots&g_{{}_{n-k}}&0&\cdots&\cdots&0\\ 0&\sigma(g_{{}_{0}})&\sigma(g_{{}_{1}})&\cdots&\sigma(g_{{}_{n-k}})&0&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&&\ddots&&\vdots\\ \vdots&&\ddots&\ddots&\ddots&&\ddots&\vdots\\ 0&\ldots&&0&\sigma^{k-1}(g_{{}_{0}})&\sigma^{k-1}(g_{{}_{1}})&\ldots&\sigma^{k-1}(g_{{}_{n-k}})\\ \end{array}\right)$
generating $C$ represents the right multiplication with $g$ in $R/Rf$, calculated with respect to the basis $1,t,\dots,t^{m-1}$, identifying $g=\sum_{i=0}^{r}g_{i}t^{i}$ with the vector $(g_{0},\dots,g_{r},0,\dots,0)$.
This generalizes [OaHA25, Theorem 2.20, 2.21] and [BagPa2025, Section 5] to the nonassociative setting. We give the proof for the convenience of the reader.
Proof.
Let $Rg/Rf$ be the left ideal in $R/Rf$ generated by $g=\sum_{i=0}^{r}g_{i}t^{i}$.
$(i)$ If $c=(c_{0},c_{1},\dots,c_{n-1})\in C$, then by definition $\sum_{i=0}^{n-1}c_{i}t^{i}\in Rg/Rf$, and so also $t\sum_{i=0}^{n-1}c_{i}t^{i}=\sum_{i=0}^{n-1}z_{1}c_{i}t^{i}\in Rg/Rf$. We compute $t\sum_{i=0}^{n-1}c_{i}t^{i}=\sum_{i=0}^{n-1}tc_{i}t^{i}=c_{0}t+c_{1}t^{2}+\dots+c_{n-2}t^{n-1}+dc_{n-1}$, hence also $c=(ac_{n-1},c_{0},c_{1}\dots,c_{n-2})\in C$.
$(ii)$ is straightforward.
(iii, iv) Since $S_{f}g$ is a left ideal, $t^{i}g\in S_{f}g$ for all $i\in\{1,\dots,n-1\}.$ It is easy to see that $g,tg,t^{2}g,\dots,t^{n-r-1}g$ are linearly independent over $K$, thus form a basis. The rows of the matrix generating $C$ are thus given by the vectors corresponding to $g,tg,t^{2}g,\dots,t^{n-r-1}g$. We obtain the generator matrix $G$ of $C$ as
(2.2) $G=\left(\begin{array}[]{cccccccc}g_{{}_{0}}&g_{{}_{1}}&\cdots&g_{{}_{n-k}}&0&\cdots&\cdots&0\\ 0&\sigma_{1}(g_{{}_{0}})&\sigma_{1}(g_{{}_{1}})&\cdots&\sigma_{1}(g_{{}_{n-k}})&0&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&&\ddots&&\vdots\\ \vdots&&\ddots&\ddots&\ddots&&\ddots&\vdots\\ 0&\ldots&&0&\sigma_{k-1}(g_{{}_{0}})&\sigma_{k-1}(g_{{}_{1}})&\ldots&\sigma_{k-1}(g_{{}_{n-k}})\\ \end{array}\right)$
where $k=n-\deg(g)$. Moreover, we just showed that the skew $(\sigma,d)$-constacyclic code $C\subset S^{n}$ corresponding to the principal left ideal $S_{f}g$ is a free left $S$-module of dimension $n-{\rm deg}(g)$. ∎
Since we assume that $\sigma\in{\rm Aut}(S)$, note that a linear code $C$ is a left skew $(f,\sigma,\delta)$-polycyclic code if and only if the corresponding polynomials form a left ideal in $S_{f}$. Then $C$ is a left skew $(f,\sigma,\delta)$-polycyclic code if and only if it is a right ideal in $\quad{}_{f}S=S^{op}[t;\sigma^{-1},-\delta\circ\sigma^{-1}]/S^{op}[t;\sigma^{-1},-\delta\circ\sigma^{-1}]f^{op}$, with $f^{op}$ being $f$ but its coefficients written on the right. (This is [BLMS2024, Corollary] formulated using ideals in our terminology.)
3. Equivalence relations for skew polycyclic codes
3.1. Background
The Hamming weight of $c=(c_{0},\dots,c_{m-1})\in S^{m}$ counts the number of nonzero components in $c$ and is defined as $w(c)=\sum_{r\not=0}n_{r}(c)$ where $n_{r}(c)=\{i\,|\,c_{i}=r\}$ counts the number of components that equals the element $r\in S$. The Hamming weight of a polynomial $p(t)\in R$ is defined as the number of nonzero coefficients of $p(t)$.
We want equivalent linear codes to have the same performance parameters, i.e. the same length $m$, dimension $k$, and minimum Hamming distance $d$. When looking at the literature, there are several slightly different ways to define two linear codes with the Hamming metric to be “equivalent”. The most general definition for commutative rings $S$ seems to be the following:
Two linear codes $C_{1}$ and $C_{2}$ over a ring $S$ of the same length $m$ are sometimes called semilinear-equivalent, if there exists an automorphism $\tau\in{\rm Aut}(S)$, $\gamma=(v,\pi)\in S^{m}\rtimes S_{m}$ such that $C_{2}=\{y\in S^{m}\,|\,y=\gamma(\tau(c))\text{ for }x\in C_{1}\}$. If $C_{1}$ and $C_{2}$ have generator matrices $G_{1}$ and $G_{2}$ this means that then they are semilinear-equivalent, if there is an invertible $N\in Gl_{n}(S)$, some $\tau\in{\rm Aut}(S)$ and a monomial matrix $M\in M_{n}(S)$ such that $G_{2}=N\tau(G_{1}M)$. If $\tau=id$ then $C_{1}$ and $C_{2}$ are called linearly-equivalent.
That means $C_{1}$ and $C_{2}$ are semilinear-equivalent if and only if one can be obtained from the other by a finite sequence of any of the following transformations, applied to an element $(c_{0},\dots,c_{m-1})$ in one code:
  1. (1)
    A permutation of its elements.
  2. (2)
    The multiplication of elements in a fixed position by a nonzero scalar in $S$.
  3. (3)
    Applying an isomorphism in ${\rm Aut}(S)$ to each component of a code vector in $S^{m}$.
This definition is motivated by the fact that the above operations all preserve the Hamming distance and the linearity of a code.
Note that the set of all $m\times m$ monomial matrices forms a group, and that every monomial matrix is the product of a permutation matrix $P$ and a diagonal matrix $diag(d_{0},\dots,d_{m-1})$ with $d_{i}\in S^{\times}$ invertible.
Chen et al. defined an equivalence relation on the invertible elements of $\mathbb{F}_{q}$ we will call Chen isometry in the following, to classify the constacyclic codes over $\mathbb{F}_{q}$ of length $m$ [CFLL2012]. Elements $a,b\in\mathbb{F}_{q}^{\times}$ in the same equivalence class define equivalent classes of constacyclic codes ${\bf C}_{t^{m}-a}$ and ${\bf C}_{t^{m}-b}$, and in turn guarantee the existence of an isometry between the two associative algebras $\mathbb{F}_{q}[t]/(t^{m}-a)$, respectively, $\mathbb{F}_{q}[t]/(t^{m}-b)$.
However, Chen isometry does not easily yield the generator polynomial of the code under the isometry. This lead to the study of Chen equivalence $\sim_{m}$ to classify the constacyclic codes over $\mathbb{F}_{q}$ [CDL2015] which is a special case of Chen isometry. It trivially allows to relate the generating polynomials of Chen equivalent constacyclic codes. Both definitions were recently generalised to skew constacyclic and (skew) polycylic codes over finite fields and rings [BBB21, CON2025, OaNA25, LM25, NevPu2025]. For instance, [LM25, Section 6] uses the monomial isometries $G_{id,\alpha}$ of degree one, when the algebras $R_{a}$ are associative.
Another approach was taken in [BLMS2024]: The classification of $(f,\sigma,\delta)$-polycyclic codes that are Hamming isometrical equivalent defines the class of $(f,\sigma,\delta)$-polycyclic codes ${\bf C}_{f}$ and the class of $(h,\sigma,\delta)$-polycyclic codes ${\bf C}_{h}$ to be Hamming isometrical equivalent, if there exists a monomial matrix $B$ such that $A_{f_{1}}B=\sigma(B)A_{f_{2}}+\delta(B)$. This definition implies that $f_{1}\sim f_{2}$ are similar skew polynomials. It is straightforward to see that if two classes of codes are Chen isometric, then they are Hamming isometrical equivalent.
3.2. Our approach
Let $S$ be a unital associative ring, $\sigma\in{\rm Aut}(S)$, $f,h\in R=S[t;\sigma,\delta]$ monic, $S_{0}$ the ring fixed under $\sigma$. We allow that $\sigma=id$ and $\delta=0$ unless stated otherwise. Let $f,h\in R$ be monic polynomials of degree $m$.
Let us define those isomorphisms of Petit algebras that corresponds to Hamming-weight preserving homomorphisms of skew polycyclic codes.
Definition 3.1.
Suppose $\tau\in\mathrm{Aut}(S)$, $\alpha\in S^{\times}$ is nonzero and $k\in\mathbb{N}$. If there exists a ring isomorphism
$G:R/Rf\to R/Rh$
defined via $G|_{S}=\tau$ and $G_{\tau,\alpha,k}(t)=\alpha t^{k}$, that is,
(3.1) $G\left(\sum_{i=0}^{m-1}a_{i}t^{i}\right)=\sum_{i=0}^{m-1}\tau(a_{i})(\alpha t^{k})^{i}.$
then we call $G$ an isometry or a monomial isomorphism of degree $k$, if we want to clarify the degree, and write $G=G_{\tau,\alpha,k}$. When $k=1$, we write $G_{\tau,\alpha}$ in place of $G_{\tau,\alpha,1}$ and call $G_{\tau,\alpha}$ an equivalence.
A ring isomorphism $G_{id,\alpha,k}:R/Rf\to R/Rh$ is called a a Chen isometry, and $G_{id,\alpha}$ a Chen-equivalence.
If $m$ is not clear from our context (which corresponds to the length of the code), we will sometimes use the terms $m$-isometry, $m$-equivalence, $m$-Chen isometry, and $m$-Chen equivalence instead.
When $\delta=0$ we require some additional notation. For any $i\in\mathbb{N}$, $\tau\in\mathrm{Aut}(S)$ and $\beta\in S^{\times}$, write
$N_{i}^{\tau}(\beta)=\prod_{j=0}^{i-1}\tau^{j}(\beta).$
Note that $\beta\tau(N_{i}^{\tau}(\beta))=N_{i+1}^{\tau}(\beta)$ and thus
(3.2) $N_{i+j}^{\tau}(\beta)=N_{i}^{\tau}(\beta)\cdot\sigma^{i}(N_{j}^{\tau}(\beta))$
for all $i,j\geq 0$. We have $N_{n}^{\sigma}(\beta)\in S_{0}$ for all $\beta\in S$ so we distinguish this case with the notation $N_{n}^{\sigma}=N_{S/S_{0}}$.
When $\delta=0$ we then can write
$G_{\tau,\alpha}\left(\sum_{i=0}^{m-1}a_{i}t^{i}\right)=\sum_{i=0}^{m-1}\tau(a_{i})N_{i}^{\sigma}(\alpha)t^{i}.$
If $S=K$ is a field, $S_{0}=K_{0}$, $K/K_{0}$ a cyclic Galois extension and $\mathrm{Gal}(K/K_{0})=\langle\sigma\rangle$, where $\sigma$ has order $m$, then $N_{n}^{\sigma}$ is simply the norm $N_{K/K_{0}}$ of the field extension $K/K_{0}$.
Up to now, only Chen isometries have been used in the literature for (skew) polycyclic codes, cf. Section 3.1. Their disadvantage is they do not include property (3) when using them to define the equivalence of (skew) polycyclic codes.
Let ${\bf C}_{f}$ be the class of skew $(f,\sigma,\delta)$-polycyclic codes and ${\bf C}_{h}$ the class of skew $(h,\sigma,\delta)$-polycyclic codes.
We know that the above defined Hamming weight preserving ring isomorphisms are the only ones there are in many cases, see Section 4. This makes this approach the most general one possible.
Definition 3.2.
The classes ${\bf C}_{f}$ and ${\bf C}_{h}$ are called
$(i)$ isometric, if there exists an isometry $G_{\tau,\alpha,k}:R/Rf\rightarrow R/Rh,$
$(ii)$ Chen isometric, if there exists an isometry $G_{id,\alpha,k}:R/Rf\rightarrow R/Rh,$
(iii) equivalent, if there exists an equivalence $G_{\tau,\alpha}:R/Rf\rightarrow R/Rh$
(iv) Chen equivalent, if there exists a Chen equivalence $G_{id,\alpha}:R/Rf\rightarrow R/Rh.$
If the length of the codes is not clear from the context, we will use the terminology $m$-isometric, $m$-Chen isometric etc. instead.
If two classes ${\bf C}_{f}$ and ${\bf C}_{h}$ are isometric or equivalent, then $f$ and $h$ have the same degree.
Equivalent, isometric, Chen equivalent or Chen isometric classes of codes ${\bf C}_{f}$ and ${\bf C}_{h}$ contain codes with the same parameters, thus have the same performance: for every skew polycyclic code $C_{1}$ in ${\bf C}_{f}$ we can find a skew polycyclic code $C_{2}$ in ${\bf C}_{h}$ that has the same length, dimension, and Hamming weight. One can be obtained from the other by a finite sequence of any of (1), (2) or (3).
We will show that in the majority of cases, the generator of a code in ${\bf C}_{f}$ can be explicitly matched with the one of a code in an equivalent class ${\bf C}_{h}$. This is because of the scarcity of monomial isomorphisms of degree $k=1$.
Remark 3.3.
If ${\bf C}_{f}$ and ${\bf C}_{h}$ are Chen equivalent, then they are also equivalent.
If ${\bf C}_{f}$ and ${\bf C}_{h}$ are Chen isometric, then they are also isometric.
If ${\bf C}_{f}$ and ${\bf C}_{h}$ are equivalent, then they are also isometric.
If ${\bf C}_{f}$ and ${\bf C}_{h}$ are Chen equivalent, then they are also equivalent and Chen isometric, hence isometric.
To motivate these notions we stress that the classifications we will achieve are tighter when employing the notions of equivalence and isometry (Section 5.2).
Inspired by [OaHA25, OuaHA25] we define equivalence relations on $(S^{\times})^{m}$ via the following equivalence relations:
Definition 3.4.
We call $(a_{0},\dots,a_{m-1})$ isometric to $(b_{0},\dots,b_{m-1})$, written
$(a_{0},\dots,a_{m-1})\cong_{m}(b_{0},\dots,b_{m-1}),$
if there exists an isometry $G_{\tau,\alpha,k}:R/Rf\rightarrow R/Rh,$ and Chen isometric, if $\tau=id$.
We call $(a_{0},\dots,a_{m-1})$ equivalent to $(b_{0},\dots,b_{m-1})$, written
$(a_{0},\dots,a_{m-1})\sim_{m}(b_{0},\dots,b_{m-1}),$
if there exists an equivalence $G_{\tau,\alpha}:R/Rf\rightarrow R/Rh$, and Chen equivalent, written
$(a_{0},\dots,a_{m-1})\sim_{Chen,m}(b_{0},\dots,b_{m-1})$
if $\tau=id$.
The above equivalence relations partition $(S^{\times})^{m}$ into different equivalence classes. It is also obvious that
$(a_{0},\dots,a_{m-1})\cong_{m}(b_{0},\dots,b_{m-1})$
if and only if ${\bf C}_{f}$ and ${\bf C}_{h}$ are isometric,
$(a_{0},\dots,a_{m-1})\cong_{Chen,m}(b_{0},\dots,b_{m-1})$
if and only if ${\bf C}_{f}$ and ${\bf C}_{h}$ are Chen isometric,
$(a_{0},\dots,a_{m-1})\sim_{m}(b_{0},\dots,b_{m-1})$
if and only if ${\bf C}_{f}$ and ${\bf C}_{h}$ are equivalent, and
$(a_{0},\dots,a_{m-1})\sim_{Chen,m}(b_{0},\dots,b_{m-1})$
if and only if ${\bf C}_{f}$ and ${\bf C}_{h}$ are Chen equivalent
Remark 3.5.
The equivalence relation $\sim_{m}$ that partitions $(S^{\times})^{m}$ into $m$-isometry classes, respectively ${\bf C}_{f}$ and ${\bf C}_{h}$ into equivalent classes of codes, is finer than the Chen equivalence and Chen isometry in [CFLL2012, OaHA25, OuaHA25], for instance see [NevPu2025, Example 5.8].
Remark 3.6.
In [OaHA25, Definition 4.1], for $K=\mathbb{F}_{q}$, two skew polynomials $f_{1},f_{2}\in K[t;\sigma]$ are called Hammming equivalent (resp. rank equivalent, if there is a $K$-vector space isomorphism $\varphi:K[t;\sigma]/K[t;\sigma]f_{1}\to K[t;\sigma]/K[t;\sigma]f_{2}$ such that
  1. (1)
    $\varphi(gh)=\varphi(g)\varphi(h)$ for all $g,h\in K[t;\sigma]/K[t;\sigma]f_{1}$
  2. (2)
    $\varphi$ preserves the Hamming distance $d_{H}(g,h)=d_{H}(\varphi(g),\varphi(h)),$ resp. the rank distance $d_{R}(g,h)=d_{H}(\varphi(g),\varphi(h))$.
Here, the rank distance is defined over arbtrary subfields of $K$.
We can now read this definition as two skew polynomials $f_{1},f_{2}\in K[t;\sigma]$ being defined as Hamming or rank equivalent, if there exist a Chen isometry $G_{id,\alpha,k}:K[t;\sigma]/K[t;\sigma]f_{1}\to K[t;\sigma]/K[t;\sigma]f_{2}$. Thus, the above mentioned - and for $f_{1}$ not two-sided still undefined -multiplication must be the one in the nonassociative Petit algebra $R/Rf_{1}$, that is viewed as algebra over some subfield of $K$ in the rank metric case.
With the results outlined in this paper, we can say that “in most cases” two skew polynomials $f_{1},f_{2}\in K[t;\sigma]$ are Hamming or rank equivalent, if there exist a Chen equivalence $G_{id,\alpha}:K[t;\sigma]/K[t;\sigma]f_{1}\to K[t;\sigma]/K[t;\sigma]f_{2}$, where we require that $\alpha\in K^{\prime\times}$ when we use the rank metric. In most cases means this is the case when $f_{1}$ is not two-sided and the degree of $f$ is greater than m1m-1, or when $f_{1}(t)=t^{m}-a$ and $f_{2}(t)=t^{m}-b$ are not two-sided. The decision in [OaHA25] to focus on the case that $\varphi(t)=\alpha t$ is thus well founded. However, to instead also consider the $K^{\prime}$-algebra isomorphisms obtained via $G_{\tau,\alpha}$ for all $\tau\in{\rm Gal}(K/K^{\prime})$ should yield tighter classifications here as well.
4. Isometries between Petit algebras $S[t;\sigma]/S[t;\sigma]f$
Let $S$ be a unital commutative associative ring and suppose $\sigma\in\mathrm{Aut}(S)$ has finite order $n$. Write $S_{0}=\{s\in S:\sigma(s)=s\}$ for the subring fixed by $\sigma$. If $\tau\in\mathrm{Aut}(S)$ is an arbitrary automorphism then we write $S_{0}^{\tau}=\{s\in S_{0}:\tau(s)=s\}$ for the common fixed subring under $\tau$ and $\sigma$. When $S=K$ is a field, we define $K_{0},K_{0}^{\tau}$ analogously.
For any $i\in\mathbb{N}$, $\tau\in\mathrm{Aut}(S)$ and $\beta\in S$, write
$N_{i}^{\tau}(\beta)=\prod_{j=0}^{i-1}\tau^{j}(\beta).$
Note that $\beta\tau(N_{i}^{\tau}(\beta))=N_{i+1}^{\tau}(\beta)$ and thus
(4.1) $N_{i+j}^{\tau}(\beta)=N_{i}^{\tau}(\beta)\cdot\sigma^{i}(N_{j}^{\tau}(\beta))$
for all $i,j\geq 0$. If $S=K$ is a field, $K/K_{0}$ a cyclic Galois extension, and $\mathrm{Gal}(K/K_{0})=\langle\sigma\rangle$, where $\sigma$ has order $n$, then $N_{n}^{\sigma}$ is simply the norm map $N_{K/K_{0}}$ corresponding to $K/K_{0}$. More generally, $N_{n}^{\sigma}(\beta)\in S_{0}$ for all $\beta\in S$ so we distinguish this case with the notation $N_{n}^{\sigma}=N_{S/S_{0}}$.
Let
$f(t)=t^{m}-\sum_{i=0}^{m-1}a_{i}t^{i},\quad h(t)=t^{m}-\sum_{i=0}^{m-1}b_{i}t^{i}\in S[t;\sigma]$
be two monic polynomials.
4.1. Equivalences
The following results were originally stated for $S_{0}$-isomorphisms for the case that $S$ is a field in [BrPum2018, Theorems 28, 29].
Theorem 4.1.
Let $\alpha\in S^{\times}$ and assume that $\tau\in{\rm Aut}(S)$ commutes with $\sigma$. Then
$G_{\tau,\alpha}:R/Rf\to R/Rh$
is an isomorphism of $S_{0}^{\tau}$-algebras if and only if
(4.2) $\tau(a_{i})=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$
for all $i\in\{0,\ldots,m-1\}$.
Proof.
Let $G=G_{\tau,\alpha}:R/Rf\to R/Rh$,
$G_{\tau,\alpha}(\sum_{i=0}^{m-1}d_{i}t^{i})=\sum_{i=0}^{m-1}\tau(d_{i})N_{i}^{\sigma}(\alpha)t^{i}$
be an isomorphism of $S_{0}^{\tau}$-algebras, $\alpha\in S^{\times}$. Then we have $G(zt^{i})=G(z)G(t)^{i}=\tau(z)(\alpha t)^{i}=\tau(z)N_{i}^{\sigma}(\alpha)t^{i}$ for all $i\in\{1,\ldots,m-1\}$ and $z\in S$.
Moreover, with $t^{m}=tt^{m-1}$, also
(4.3) $G(t^{m})=G\Big{(}\sum_{i=0}^{m-1}a_{i}t^{i}\Big{)}=\sum_{i=0}^{m-1}G(a_{i})G(t)^{i}=\sum_{i=0}^{m-1}\tau(a_{i})\big{(}\prod_{l=0}^{i-1}\sigma^{l}(\alpha)\big{)}t^{i}=\sum_{i=0}^{m-1}\tau(a_{i})N_{i}^{\sigma}(\alpha)t^{i}$
and $G(tt^{m-1})=G(t)G(t^{m-1})=G(t)G(t)^{m-1}$, i.e.
(4.4) $G(t)^{m}=G(t)G(t)^{m-1}=\alpha\sigma(\alpha)\cdots\sigma^{m-1}(\alpha)t^{m}=N^{\sigma}_{m}(\alpha)\sum_{i=0}^{m-1}b_{i}t^{i}.$
Comparing (4.3) and (4.4) gives $\tau(a_{i})=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$ for all $i\in\{0,\ldots,m-1\}$. Thus Equation (4.2) holds for all $i\in\{0,\ldots,m-1\}$.
Conversely, if Equation (4.2) holds for all $i\in\{0,\ldots,m-1\}$ and some invertible $\alpha\in S$, then consider the isomorphism $G_{\tau,\alpha}:R\rightarrow R$. By Lemma 2.5, this isomorphism canonically induces an algebra isomorphism between the Petit algebras $R/Rf$ and $R/RG_{\tau,\alpha}(f)$. Since
$G_{\tau,\alpha}(f)=N_{m}^{\sigma}(\alpha)t^{m}-\sum_{i=0}^{m-1}\tau(a_{i})N_{i}^{\sigma}(\alpha)t^{i}$
$N_{m}^{\sigma}(\alpha)(t^{m}-\sum_{i=0}^{m-1}\tau(a_{i})N_{m}^{\sigma}(\alpha^{-1})N_{i}^{\sigma}(\alpha)t^{i})=N_{m}^{\sigma}(\alpha)(t^{m}-\sum_{i=0}^{m-1}\tau(a_{i})N_{m-i}^{\sigma}(\sigma^{i}(\alpha^{-1}))t^{i})$
$N_{m}^{\sigma}(\alpha)(t^{m}-\sum_{i=0}^{m-1}b_{i}t^{i})$
we obtain again $\tau(a_{i})=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$ for all $i\in\{0,\ldots,m-1\}$ and thus the assertion. ∎
When $K[t;\sigma]/K[t;\sigma]f$ and $K[t;\sigma]/K[t;\sigma]h$ are proper nonassociative, and $\mathrm{Aut}(K)$ is abelian, we can say even more.
Theorem 4.2.
Let $n\geq m-1$ and assume that $f$ and $h$ do not generate two-sided ideals in $K[t;\sigma]$. Suppose $\mathrm{Aut}(K)$ is abelian. For any subfield $F\subset K_{0}$, the only $F$-algebra isomorphisms
$G:K[t;\sigma]/K[t;\sigma]f\to K[t;\sigma]/K[t;\sigma]h$
are of the form $G=G_{\tau,\alpha}$ such that $\tau\in\mathrm{Aut}(K)$, $K_{0}^{\tau}\subset F$, and $\alpha\in K^{\times}$ satisfying Equation 4.2
Proof.
Let $G:K[t;\sigma]/K[t;\sigma]f\to K[t;\sigma]/K[t;\sigma]h$ be an isomorphism of $F$-algebras. Since $K[t;\sigma]/K[t;\sigma]f$ is not associative, its left nucleus is $K$. Since any isomorphism preserves the left nucleus, $G(K)=K$ and so $G|_{K}=\tau$ for some $\tau\in\text{Aut}_{F}(K)$. Suppose $G(t)=\sum_{i=0}^{m-1}k_{i}t^{i}$ for some $k_{i}\in K$. Then we have
(4.5) $G(tz)=G(t)G(z)=(\sum_{i=0}^{m-1}k_{i}t^{i})\tau(z)=\sum_{i=0}^{m-1}k_{i}\sigma^{i}(\tau(z))t^{i}$
and
(4.6) $G(tz)=G(\sigma(z)t)=\sum_{i=0}^{m-1}\tau(\sigma(z))k_{i}t^{i}$
for all $z\in K$. Comparing the coefficients of $t^{i}$ in (4.5) and (4.6) we obtain
(4.7) $k_{i}\sigma^{i}(\tau(z))=k_{i}\tau(\sigma^{i}(z))=\tau(\sigma(z))k_{i}=k_{i}\tau(\sigma(z))$
for all $i\in\{0,\ldots,m-1\}$ and all $z\in K$. This implies $k_{i}(\tau\big{(}\sigma^{i}(z)-\sigma(z)\big{)})=0$ for all $i\in\{0,\ldots,m-1\}$ and all $z\in K$ since $\sigma$ and $\tau$ commute, i.e.
(4.8) $k_{i}=0\text{ or }\sigma^{i}(z)=\sigma(z)$
for all $i\in\{0,\ldots,m-1\}$ and all $z\in K$.
Since $\sigma$ has order $n\geq m-1$, which means $\sigma^{i}\not=\sigma$ for all $i\in\{0,\ldots,m-1\}$, $i\not=1$, (4.8) implies $k_{i}=0$ for all $i\in\{0,\ldots,m-1\}$, $i\not=1$. Therefore $G(t)=\alpha t$ for some $\alpha\in K^{\times}$. As in the proof of Theorem 4.1, these maps are indeed well defined algebra isomorphisms and must be the only one which implies the assertion. ∎
Proposition 4.3.
There exists an algebra isomorphism $G_{\tau,\alpha}:S_{f}\rightarrow S_{h}$ if and only if there exist algebra isomorphisms $G_{\tau,\sigma^{i}(\alpha)}:S_{t^{m-i}-a_{i}}\rightarrow S_{t^{m-i}-b_{i}}$ for all $i\in\{0,\ldots,m-1\}$.
Proof.
We know that $G_{\tau,\alpha}$ is an algebra isomorphism if and only if
$\tau(a_{i})=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$
for all $i\in\{0,\ldots,m-1\}$. Put $\alpha_{i}=\sigma^{i}(\alpha)$, then the above equations are the same as saying that there exist $\alpha_{i}\in K^{\times}$ such that
$\tau(a_{i})=N_{m-i}^{\sigma}(\beta_{i})b_{i}$
which is the same as saying that the algebras $S_{t^{m-i}-a_{i}}$ and $S_{t^{m-i}-b_{i}}$ are isomorphic with the isomorphisms given by $G_{\tau,\alpha_{i}}$. ∎
Theorem 4.4.
Suppose $\tau\in{\rm Aut}(S)$ commutes with $\sigma$ and define
$f^{\tau}(t)=t^{m}-\sum_{i=0}^{m-1}\tau(a_{i})t^{i}.$
Then
$G_{\tau,\alpha}:R/Rf\to R/Rh$
is an isomorphism if and only if
$G_{id,\alpha}:R/Rf^{\tau}\to R/Rh$
is an isomorphism.
Proof.
By Theorem 4.1, $G_{\tau,\alpha}:R/Rf\to R/Rh$ is an isomorphism of $S_{0}^{\tau}$-algebras if and only if $\tau(a_{i})=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$ for all $i\in\{0,\ldots,m-1\}$. But this is equivalent to $G_{\tau,\alpha}:R/Rf^{\tau}\to R/Rh$ being an isomorphism. ∎
4.2. Isometries
It may happen that $G(t)$ is a polynomial of degree $k>1$ for some isomorphism $G$ between two nonassociative algebras $R/Rf$ and $R/Rh$. We are only interested in Hamming weight preserving isomorphism, so we are concerned with the cases when $G(t)=\alpha t^{k}$ for some $k>1$ and $\alpha\in S^{\times}$ invertible, i.e. when $G=G_{\tau,\alpha,k}$ is monomial of degree $k$.
Note that when $S$ is a field and $\sigma\not=id$, that case can only ever occur when n<m1n<m-1. Isometry classes are the hardest to understand. We collect some partial results here to make our point.
Lemma 4.5.
Let $\sigma\not=id$. Let $\tau\in\mathrm{Aut}(S)$ commute with $\sigma$, $k\in\mathbb{N}$, 1<k<m11<k<m-1, and $\alpha\in S^{\times}$. If the map $G$ denoted $G_{\tau,\alpha,k}$ which is defined via $G(c)=\tau(c)$ for all $c\in S$ and $G(t)=\alpha t^{k}$ defines an isomorphism between $R/Rf$ and $R/Rh$ then $k\equiv 1\mod m$.
Proof.
Take $G=G_{\tau,\alpha,k}:S[t;\sigma]\to S[t;\sigma]$ $G(\sum_{i=0}^{\ell}a_{i}t^{i})=\sum_{i=0}^{\ell}\tau(a_{i})N_{i}^{\sigma}(\alpha)t^{ik},$ for any $\ell\in\mathbb{N}$ and $a_{i}\in S$. The map $G=G_{\tau,\alpha,k}$ is $S_{0}^{\tau}$-linear and multiplicative, therefore we have $G(t)G(c)=G(\sigma(c))G(t)$ for all $c\in S$. We compute
$G(t)G(c)=\alpha t^{k}\tau(c)=\alpha\sigma^{k}(\tau(c))t^{k}$
and
$G(\sigma(c))G(t)=\tau(\sigma(c))\alpha t^{k}.$
Since $\alpha\in S^{\times}$ and $\sigma$ and $\tau$ commute, we must have $\sigma=\sigma^{k}$, or $k\equiv 1\mod m$. ∎
Lemma 4.6.
Let $\sigma\not=id$. Let $\tau\in\mathrm{Aut}(S)$ commute with $\sigma$, $k\in\mathbb{N}$, 1<k<m11<k<m-1, and $\alpha\in S^{\times}$. If the map $G_{\tau,\alpha,k}$ which is defined via $G(c)=\tau(c)$ for all $c\in S$ and $G(t)=\alpha t^{k}$ defines an isomorphism between $R/Rf$ and $R/Rh$ then
$N^{\sigma^{k}}_{m}(\alpha)(\sum_{i=0}^{m-1}b_{i}t^{i})^{k}=\sum_{i=0}^{m-1}\tau(a_{i})N_{i}^{\sigma^{k}}(\alpha)t^{ki}\,mod_{r}\,f.$
Proof.
Since for all i<m1i<m-1 the powers $t^{i}$ are power associative, so must be their images, that is the powers $G(t)^{i}$ of $G(t)$ are power associative, for all $i\leq m-1$. Hence we have
$G(\sum_{i=0}^{m-1}a_{i}t^{i})=\sum_{i=0}^{m-1}\tau(a_{i})G(t)^{i}=\sum_{i=0}^{m-1}\tau(a_{i})N_{i}^{\sigma^{k}}(\alpha)t^{ki}\,mod_{r}\,f.$
Since $G$ is a homomorphism and with $t^{m}=tt^{m-1}$ this implies that
$G(t^{m})=\sum_{i=0}^{m-1}\tau(a_{i})G(t)^{i}=\sum_{i=0}^{m-1}\tau(a_{i})N_{i}^{\sigma^{k}}(\alpha)t^{ki}\,mod_{r}\,f$
$N_{m}^{\sigma^{k}}(\alpha)(\sum_{i=0}^{m-1}a_{i}t^{i})^{k}=\sum_{i=0}^{m-1}\tau(a_{i})N_{i}^{\sigma^{k}}(\alpha)t^{ki}.$
Now $G(tt^{m-1})=G(t)G(t^{m-1})=G(t)G(t)^{m-1}$, i.e.
$G(t)^{m}=(\alpha t^{k})(\alpha t^{k})^{m-1}=N_{m}^{\sigma^{k}}(\alpha)(t^{k})^{m}=N_{m}^{\sigma^{k}}(\alpha)(t^{m})^{k}=N^{\sigma^{k}}_{m}(\alpha)(\sum_{i=0}^{m-1}b_{i}t^{i})^{k}.$
(Note we are computing in the associative skew polynomial ring here.) Comparing these two equations yields the assertion. ∎
Proposition 4.7.
Let $\sigma\not=id$. Let $G:R/Rf\rightarrow R/Rh$ be an isomorphism such that $G|_{K}=\tau$ for some $\tau\in{\rm Aut}(S)$ and $G(t)=\alpha t^{k}$ for some $k>1$. If any of the following conditions fail, then $G$ is not well defined:
  1. (1)
    $k\equiv 1\mod n$ $1<k<m$;
  2. (2)
    $\gcd(k,m)=1$;
  3. (3)
    if $\ell$ is the inverse of $k$ mod $m$, then $\ell=1+jn$ for some $1<\ell<m$;
  4. (4)
    $N^{\sigma^{k}}_{m}(\alpha)(\sum_{i=0}^{m-1}b_{i}t^{i})^{k}=\sum_{i=0}^{m-1}\tau(a_{i})N_{i}^{\sigma^{k}}(\alpha)t^{ki}\,mod_{r}\,f.$
Proof.
The first condition is a consequence of the requirement that $G(ta)=G(\sigma(a)t)$, cf. Lemma 4.5.
If $G$ is an isomorphism, then $t=G(r)$ for some $r\in R$. Since $G$ preserves Hamming weight, $r=st^{\ell}$ for some $\ell$ and $s\in K^{\times}$. If $G$ is a homomorphism, then
$G(st^{\ell})=\tau(s)G(t)^{\ell}\in Kt^{k\ell}\subset R^{\prime}.$
That is, we must have $k\ell\equiv 1\mod m$. In particular, $k$ is relatively prime to $m$.
If $G$ is an isomorphism, then its inverse must also be an isomorphism. Therefore by the first condition, we must have $\ell\equiv 1\mod n$. This settles (2) and (3).
Condition (4) is proved in Lemma 4.6. ∎
Corollary 4.8.
Let $G:S[t;\sigma]/S[t;\sigma](t^{m}-a)\to S[t;\sigma]/S[t;\sigma](t^{m}-b)$ be an isomorphism such that $G|_{S}=\tau$ for some $\tau\in{\rm Aut}(S)$ and $G(t)=\alpha t^{k}$ for some $k>1$. Then $N^{\sigma^{k}}_{m}(\alpha)b^{k}=\tau(a)$.
Proof.
Here, $N^{\sigma^{k}}_{m}(\alpha)(\sum_{i=0}^{m-1}b_{i}t^{i})^{k}=\sum_{i=0}^{m-1}\tau(a_{i})N_{i}^{\sigma^{k}}(\alpha)t^{ki}\,mod_{r}\,f$ becomes $N^{\sigma^{k}}_{m}(\alpha)b^{k}=\tau(a)$. The assumption that $\sigma\not=id$ is not required here. ∎
In the special case when $S[t;\sigma]/S[t;\sigma](t^{m}-a)$ is associative we see an interplay between the two notions of isometry and equivalence.
Theorem 4.9.
[NevPum2025, Theorem 5.9] Suppose that $a,b\in S_{0}^{\times}$ and $n|m$. Assume that $\tau$ commutes with $\sigma$. Then the following statements are equivalent for any integer $1\leq k<m$:
  1. $(i)$
    $S[t;\sigma]/S[t;\sigma](t^{m}-a)$ and $S[t;\sigma]/S[t;\sigma](t^{m}-b)$ are isometric via $G_{\tau,\alpha,k}$;
  2. $(ii)$
    $S[t;\sigma]/S[t;\sigma](t^{m}-b^{k})$ and $S[t;\sigma]/S[t;\sigma](t^{m}-a)$ are equivalent via $G_{\tau,\alpha}$, where $k$ satisfies $k\equiv 1\mod n$ and $\gcd(k,m)=1$.
  3. $(iii)$
    $\tau(a)=N_{m}^{\sigma}(\alpha)b^{k}$ where $k\equiv 1\mod n$ and $\gcd(k,m)=1$.
Corollary 4.10.
[NevPu2025, Corollary 5.10] Suppose $K=\mathbb{F}_{p^{r}}$ and $\sigma(x)=x^{p^{s}}$ so that $n=r/s$. Suppose that $\mathbb{F}_{p^{r}}[t;\sigma]/\mathbb{F}_{p^{r}}[t;\sigma](t^{m}-a)$ is associative, that is, $a\in\mathbb{F}_{p^{s}}^{\times}$ and $n|m$. Then the number of different Chen equivalence classes of skew $(\sigma,a)$-constacyclic codes over $\mathbb{F}_{p^{r}}$ equals the number of different Chen isometry classes of skew $(\sigma,a)$-constacyclic codes.
For finite rings we can count the number of Chen isometry classes of skew constacyclic codes for classes of codes with a nonassociative ambient algebra.
Theorem 4.11.
Let $a\in S^{\times}$ and assume that $S$ is a finite ring. Then the number of distinct Chen isometry classes of families of skew $(\sigma,a)$-constacyclic codes arising from nonassociative algebras $S[t,\sigma]/S[t,\sigma](t^{m}-a)$ is
$N=\begin{cases}|S^{\times}|/|N_{m}^{\sigma}(S^{\times})|&\text{if $n\nmid m$; and}\\ |S^{\times}|/|N_{m}^{\sigma}(S^{\times})|-|S_{0}^{\times}|/|N_{m}^{\sigma}(S^{\times})|&\text{if $n|m$.}\end{cases}$
There are additionally $|S_{0}^{\times}|/|N_{m}^{\sigma}(S^{\times})|$ different Chen equivalence classes (and thus at most this many Chen isometry classes) of families of skew $(\sigma,a)$-constacyclic codes arising from associative algebras $S[t,\sigma]/S[t,\sigma](t^{m}-a)$, that is, for which $n\mid m$ and $a\in S_{0}^{\times}$.
Proof.
For any $a,b\in S^{\times}$, the class of skew $(\sigma,a)$-constacyclic codes is equivalent to the class of skew $(\sigma,b)$-constacyclic codes if and only if $b\in aN_{m}^{\sigma}(\alpha)$ for some $\alpha\in S^{\times}$, that is, if and only if $a$ and $b$ lie in the same coset of the subgroup $N_{n}^{\sigma}(S^{\times})$.
First suppose that $n\nmid m$, so that all of the algebras $S[t;\sigma]/S[t;\sigma](t^{m}-a)$ are not associative. In this case, every coset $aN_{m}^{\sigma}(S^{\times})$ corresponds to a class of skew $(\sigma,a)$-constacyclic codes such that $S[t;\sigma]/S[t;\sigma](t^{m}-a)$ is not associative, and there are $|S^{\times}|/|N_{m}^{\sigma}(S^{\times})|$ such classes.
Now suppose $n|m$, so that we are in the nonassociative case if and only if $a,b\in S\setminus S_{0}$. Then $N_{m}^{\sigma}(S^{\times})\subset S_{0}$, so that every coset of $N_{m}^{\sigma}(S^{\times})$ is either entirely contained in $S_{0}$ or entirely disjoint from it (so corresponds to a nonassociative Petit algebra).
The total number of cosets contained in $S_{0}$, and thus corresponding to associative algebras $S[t;\sigma]/S[t;\sigma](t^{m}-a)$ with $a\in S_{0}^{\times}$, is equal to
$|S_{0}^{\times}|/|N_{m}^{\sigma}(S^{\times})|.$
In that case, we have shown that the notions of Chen-isometry and Chen-equivalence need not coincide. Thus these Chen-isometry classes may be bigger (thus fewer in number), yielding the last sentence of the theorem.
Finally, subtracting these $|S_{0}^{\times}|/|N_{m}^{\sigma}(S^{\times})|$ cosets from the total number of cosets gives the number of Chen-isometry classes for nonassociative algebras in the case that $n|m$, as required. ∎
4.3. A closer look at Equation (4.2)
We now derive some conditions that help us check when two skew polycyclic codes cannot be equivalent employing Equation (4.2).
Proposition 4.12.
Let $R=S[t;\sigma]$. If $R/Rf\cong R/Rh$ via some $G_{\tau,\alpha}$ then $a_{i}=0$ if and only if $b_{i}=0$, for all $i\in\{0,\ldots,m-1\}$.
Proof.
If $R/Rf\cong R/Rh$ via some $G_{\tau,\alpha}$, then by Theorem 4.2, there exists $j\in\left\{0,\ldots,n-1\right\}$ and $\alpha\in S^{\times}$ such that $\tau(a_{i})=(\prod_{l=i}^{m-1}\sigma^{l}(\alpha))b_{i}$ for all $i\in\{0,\ldots,m-1\}$. This implies $a_{i}=0$ if and only if $b_{i}=0$, for all $i\in\{0,\ldots,m-1\}$. ∎
Proposition 4.13.
Let $\tau\in\mathrm{Aut}(S)$ commute with $\sigma$ and set $N_{S/S_{0}}=N_{n}^{\sigma}$. Let
$\tau(a_{i})=(\prod_{l=i}^{m-1}\sigma^{l}(\alpha))b_{i}$
for all $i\in\{0,\dots,m-1\}$. Suppose there exists $i_{0}\in\{0,\dots,m-1\}$ such that $N_{S/S_{0}}(\tau(a_{i_{0}}))=N_{S/S_{0}}(b_{i_{0}})$ is invertible, and that there is no nontrivial $(m-i_{0})$th root of unity in $S_{0}$. Then $N_{S/S_{0}}(\alpha)=1$.
Proof.
Since $\tau(a_{i})=(\prod_{l=i}^{m-1}\sigma^{l}(\alpha))b_{i}$ for all $i\in\{0,\ldots,m-1\}$ we get
$N_{S/S_{0}}(\tau(a_{i}))=\prod_{l=i}^{m-1}N_{S/S_{0}}(\sigma^{l}(\alpha))N_{S/S_{0}}(b_{i})=N_{S/S_{0}}(\alpha)^{m-i}N_{S/S_{0}}(b_{i})$
for all $i\in\{0,\dots,m-1\}$ by applying $N_{S/S_{0}}$ to both sides. For all $i\in\{0,\dots,m-1\}$ such that $N_{S/S_{0}}(\tau(a_{i}))=N_{S/S_{0}}(b_{i})$ is invertible, this yields $1=N_{S/S_{0}}(\alpha)^{m-i}$, therefore $N_{S/S_{0}}(\alpha)\in S_{0}^{\times}$ must be an (mi)(m-i)th root of unity. This yields the assertion. ∎
Corollary 4.14.
We assume the setting of Proposition 4.13.
$(i)$ Suppose that $N_{S/S_{0}}(\tau(a_{0}))=N_{S/S_{0}}(b_{0})$ is invertible, and that there is no nontrivial $m$th root of unity in $S_{0}$. Then $N_{S/S_{0}}(\alpha)=1$.
$(ii)$ Suppose that $N_{S/S_{0}}(\tau(a_{m-1}))=N_{S/S_{0}}(b_{m-1})\not=0$. Then $N_{S/S_{0}}(\alpha)=1$.
Lemma 4.15.
We assume the setting of Proposition 4.13.
$(i)$ If $b_{m-1}=\tau(a_{m-1})$ is invertible, then $\alpha=1$.
$(ii)$ If there is some $i_{0}\in\{0,\ldots,m-1\}$ such that $\tau(a_{i_{0}})\not=b_{i_{0}}$ then $\alpha\not=1$.
Proof.
$(i)$ If $b_{m-1}=\tau(a_{m-1})$ is invertible, then $\tau(a_{m-1})=\sigma^{m-1}(\alpha)a_{m-1}$ means $\sigma^{m-1}(\alpha)=1$, i.e. $\alpha=1$.
$(ii)$ If $\alpha=1$ then $\tau(a_{i})=b_{i}$ for all $i\in\{0,\ldots,m-1\}$. ∎
As some straightforward applications of our results, we obtain some useful equivalences which have been already observed in various literature, but to our knowledge only when $S$ is a field.
Corollary 4.16.
Let $a,b\in S^{\times}$.
$(i)$ The classes of skew $(\sigma,a)$-constacyclic codes and skew $\sigma$-cyclic codes of length $m$ over $S$ are equivalent if and only if $a\in N_{m}^{\sigma}(S^{\times}).$
Equivalence and Chen equivalence coincide. Isometry and equivalence coincide when $m$ does not divide $n$.
$(ii)$ The classes of skew $(\sigma,a)$-constacyclic codes and skew $\sigma$-negacyclic codes of length $m$ are equivalent if and only if $a\in N_{m}^{\sigma}(S^{\times}).$
Equivalence and Chen equivalence coincide. Isometry and equivalence coincide when $m$ does not divide $n$.
(iii) The $(\sigma,a)$-constacyclic codes of length $m$ are equivalent to the cyclic codes of length $m$ if and only if if there exists $\alpha\in S^{\times}$ such that $\alpha^{m}a=1$.
We allow that $\sigma=id$ in which case the statements hold for constacyclic codes.
5. Code equivalence and isometry
5.1. Equivalence, and when it coincides with isometry
Let
$f(t)=t^{m}-\sum_{i=0}^{m-1}a_{i}t^{i},\quad h(t)=t^{m}-\sum_{i=0}^{m-1}b_{i}t^{i}\in S[t;\sigma]$
be two monic polynomials.
Theorem 5.1.
Let $\alpha\in S^{\times}$ and assume that $\tau\in{\rm Aut}(S)$ commutes with $\sigma$. Then two classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes of length $m$ over $S$ are equivalent if and only if there exists $\tau\in{\rm Aut}(S)$ that commutes with $\sigma$ and $\alpha\in S^{\times}$ such that
$\tau(a_{i})=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$
for all $i\in\{0,\ldots,m-1\}$ (resp., Chen-equivalent if and only if above equation is true for $\tau=id$).
This is a corollary of Theorem 4.1 and generalizes parts of [OaHA25, Theorem 4.20] to our notion of equivalence and arbitrary rings.
Inspired by [OaHA25, Theorem 4.20], we can rephrase this and say that ${\bf C}_{f}$ and ${\bf C}_{h}$ are equivalent if and only if there exists $\alpha\in S^{\times}$ and $\tau\in{\rm Aut}(S)$ that commutes with $\sigma$ such that
$\left(\tau(a_{0}),\dots,\tau(a_{i}),\dots,\tau(a_{m-1})\right)$
$\left(N_{n}^{\sigma}(\alpha),\dots,N_{m-i}^{\sigma}(\sigma^{i}(\alpha)),\dots,N_{m-1}^{\sigma}(\sigma^{m-1}(\alpha))\right)\left(b_{0},\dots,b_{i},\dots,b_{m-1}\right)$
with componentwise multiplication.
Recall that each monic right divisor $g(t)=\sum_{i=0}^{r}g_{i}t^{i}$ of $f$ generates a principal left ideal in $R/Rf=S[t;\sigma]/S[t;\sigma]f$ which is in one-one correspondence with the generator matrix of a skew polycyclic code in ${\bf C}_{f}$. For equivalent codes, the image
$G_{\tau,\alpha}(g(t))=\sum_{i=0}^{r}\tau(g_{i})(kt)^{i}$
$g$ of $f$ under the isometry $G_{\tau,\alpha}:R/Rff\rightarrow R/Rh$ yields the generator matrix of a skew polycyclic code in ${\bf C}_{h}$ that corresponds to the left ideal in $R/Rh$ generated by $G_{\tau,\alpha}(g)$ and is straightforward to compute.
When we only consider skew constacyclic codes then we can say more.
Theorem 5.2.
[NevPu2025, Corollary 4.5]
Then the classes of skew $(\sigma,a)$-constacyclic and skew $(\sigma,b)$-constacyclic codes of length $m$ over $S$ are equivalent if and only if there exists some $\tau\in{\rm Aut}(S)$ that commutes with $\sigma$ and some $\alpha\in S^{\times}$ such that
(5.1) $\tau(a)=N_{m}^{\sigma}(\alpha)b,$
and are are Chen equivalent if and only if $\tau=id$.
If $a,b\in S\setminus S_{0}$, or $a,b\in S_{0}^{\times}$ and $n\nmid m$ then equivalence and isometry coincide.
Proposition 5.3.
Set
$f^{\tau}(t)=t^{m}-\sum_{i=0}^{m-1}\tau(a_{i})t^{i}.$
Then the following statements are equivalent:
$(i)$ The classes of skew $(f,\sigma)$-polycyclic codes and skew $(h,\sigma)$-polycyclic codes of length $m$ over $S$ are equivalent.
$(ii)$ The classes of skew $(f^{\tau},\sigma)$-polycyclic and $(h,\sigma)$-polycyclic codes of length $m$ over $S$ are Chen-equivalent for all $\tau\in{\rm Aut}(S)$ that commute with $\sigma$.
This is a corollary from Theorem 4.4.
Proof.
$(i)\Leftrightarrow(ii)$ There is a Chen isometry $G_{id,\alpha}:R/f\to R/Rh$ if and only if $a_{i}=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$ for all $i\in\{0,\ldots,m-1\}$. This is the same as $\tau^{-1}(\tau(a_{i}))=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$ for all $i\in\{0,\ldots,m-1\}$ and any automorphism $\tau$. For any $\tau$ that commutes with $\sigma$, these equalities are equivalent to the existence of an isometry $G_{\tau^{-1},\alpha}:R/f^{\tau}\to R/Rh$.
Thus $G_{id,\alpha}:R/f\to R/Rh$ if and only if $G_{\tau^{-1},\alpha}:R/f^{\tau}\to R/Rh$ for all $\tau$ that commute with $\sigma$. ∎
Moreover, $(ii)$ implies that $R/Rf$ and $R/Rf^{\tau}$ and are equivalent ambient algebras for all $\tau$ that commute with $\sigma$. We can rephrase this: if ${\bf C}_{f}$ and ${\bf C}_{h}$ are Chen-isometric then ${\bf C}_{f^{\tau}}$ and ${\bf C}_{f}$ are equivalent for all $\tau$ that commute with $\sigma$.
Proposition 5.4.
Suppose $\tau\in{\rm Aut}(S)$ commutes with $\sigma$. Then the following are equivalent:
$(i)$ The classes of skew $(f,\sigma)$-polycyclic codes and skew $(h,\sigma)$-polycyclic codes of length $m$ over $S$ are Chen-equivalent.
$(ii)$ The classes of skew $(f^{\tau},\sigma)$-polycyclic and $(h,\sigma)$-polycyclic codes of length $m$ over $S$ are equivalent.
This follows directly from Proposition 5.3.
Corollary 5.5.
Suppose that $\tau(a_{i})=a_{i}$ for all $i$. Then the following are equivalent:
$(i)$ The classes of skew $(f,\sigma)$-polycyclic and skew $(h,\sigma)$-polycyclic codes of length $m$ over $S$ are equivalent.
$(ii)$ The classes of skew $(f,\sigma)$-polycyclic and skew $(h,\sigma)$-polycyclic codes of length $m$ over $S$ are Chen-equivalent.
We now construct equivalence classes for a class of skew polycyclic codes. For a given $h(t)=t^{m}-\sum_{i=0}^{m-1}b_{i}t^{i}\in S[t;\sigma]$ define
$h_{\tau,\alpha}(t)=t^{m}-\sum_{i=0}^{m-1}N_{m-i}^{\sigma}(\sigma^{i}(\tau(\alpha)))\tau(b_{i})t^{i}\in S[t;\sigma].$
Theorem 5.6.
The equivalence class ${\bf C}_{h}]$ of classes of skew polycyclic codes of length $m$ over $S$ that are equivalent to ${\bf C}_{h}$ (respectively, Chen equivalent) is given by
${\bf C}_{h}]=\{{\bf C}_{h_{\tau,\alpha}}\,|\,\alpha\in S^{\times},\tau\in{\rm Aut}(S)\text{ such that }\tau\text{ commutes with }\sigma\}$
and
${\bf C}_{h}]_{Chen}=\{{\bf C}_{h_{id,\alpha}}\,|\,\alpha\in S^{\times}\}.$
The equivalences between $R/Rh_{\tau,\alpha}$ and $R/Rh$ are given by $G_{\tau^{-1},\tau(\alpha)}$, respectively between ${\bf C}_{h_{id,\alpha}}$ and ${\bf C}_{h}$ by $G_{id,\alpha}$.
Proof.
We start with $h(t)=t^{m}-\sum_{i=0}^{m-1}b_{i}t^{i}\in S[t;\sigma]$. Take any $\alpha\in S^{\times}$. We now define new skew polynomials that by definition will be equivalent to $h$. First define
$h_{i}:=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$
for all $i\in\{0,\ldots,m-1\}$ and
$h_{id,\alpha}(t)=t^{m}-\sum_{i=0}^{m-1}h_{i}t^{i}=t^{m}-\sum_{i=0}^{m-1}N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}t^{i}.$
Then there exists a Chen equivalence $G_{id,\alpha}$ between $R/Rh_{id,\alpha}$ and $R/Rh$ by definition of $h_{id,\alpha}$. Define
$h_{i}^{\prime}:=\tau(h_{i})=N_{m-i}^{\sigma}(\sigma^{i}(\tau(\alpha)))\tau(b_{i}).$
Note that we can write
$\tau^{-1}(h_{i}^{\prime})=\tau^{-1}(\tau(h_{i}))=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$
for all $i\in\{0,\ldots,m-1\}$.
For any $\tau\in{\rm Aut}(S)$ commuting with $\sigma$, put
$h_{\tau,\alpha}=t^{m}-\sum_{i=0}^{m-1}h_{i}^{\prime}t^{i}=t^{m}-\sum_{i=0}^{m-1}N_{m-i}^{\sigma}(\sigma^{i}(\tau(\alpha)))\tau(b_{i})t^{i}.$
Thus there exists an equivalence $G_{\tau^{-1},\tau(\alpha)}$ between $R/Rh_{\tau,\alpha}$ and $R/Rh$ by definition of $h_{id,\alpha}$. ∎
Proposition 4.3 implies an intricate connection between classes of equivalent $(\sigma,f)$-polycyclic and equivalent $(\sigma,a)$-constacyclic codes over commutative rings $S$ (the case that $\sigma=id$):
Proposition 5.7.
Two classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes over $S$ of length $m$ are equivalent via $G_{\tau,\alpha}$ if and only if for all $i\in\{0,\ldots,m-1\}$ where $a_{i}\not=0$, the class of $(\sigma,a_{i})$-constacyclic codes is (mi)(m-i)-equivalent to the class of $(\sigma,b_{i})$-constacyclic codes via $G_{\tau,\sigma^{i}(\alpha)}$.
For skew constacyclic codes more is known.
Theorem 5.8.
[NevPu2025] Suppose that all $\tau\in{\rm Aut}(K)$ commute with $\sigma$. Suppose $m\nmid n$, or that one of $a$ or $b$ is not in $F$. Then two classes of skew $(\sigma,a)$-constacyclic codes and skew $(\sigma,b)$-constacyclic codes of length $n$ are isometric if and only if they are equivalent.
Theorem 5.9.
[NevPu2025] Suppose that $a,b\in F^{\times}$ and $m|n$. Assume that $\tau$ commutes with $\sigma$. Then the following statements are equivalent:
(1) $K[t;\sigma]/K[t;\sigma](t^{n}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{n}-b)$ are isometric via $G_{\tau,\alpha,k}$;
(2) $K[t;\sigma]/K[t;\sigma](t^{n}-b^{k})$ and $K[t;\sigma]/K[t;\sigma](t^{n}-a)$ are equivalent via $G_{\tau,\alpha}$, where $k$ satisfies $k\equiv 1\mod m$ and $\gcd(k,n)=1$;
(3) $\tau(a)=N_{n}^{\sigma}(\alpha)b^{k}$.
Corollary 5.10.
Let $a,b\in S^{\times}$.
$(i)$ The classes of skew $(\sigma,a)$-constacyclic codes and of skew $\sigma$-cyclic codes of length $m$ are equivalent if and only if $a\in N_{m}^{\sigma}(S^{\times}).$
$(ii)$ The classes of skew $(\sigma,a)$-constacyclic codes and skew $\sigma$-negacyclic codes of length $m$ are equivalent if and only if $a\in N_{m}^{\sigma}(S^{\times}).$
In both $(i)$ and $(ii)$, isometry and equivalence coincide when $m$ does not divide $n$.
5.2. Skew polycyclic codes over finite fields
We can now get some straightforward results for skew polycyclic codes over finite fields. More detailed results will be obtained in a future paper.
Let $K=\mathbb{F}_{q}=\mathbb{F}_{p^{r}}$ be a finite field. Let $\xi$ be a primitive element of $K$; it has order $p^{r}-1$. Let $\sigma(x)=x^{p^{s}}$ and $K_{0}=\mathbb{F}_{p^{s}}$ be the fixed field of $\sigma$. The order of $\sigma$ is $n=r/s$. Define
$m]_{s}=\frac{p^{sm}-1}{p^{s}-1}=p^{s(m-1)}+p^{s(m-2)}+\cdots+p^{s}+1.$
Then $N_{m}^{\sigma}(x)=x^{[m]_{s}}$ for all $x\in\mathbb{F}_{p^{r}}$. The subgroup $N_{m}^{\sigma}(\mathbb{F}_{p^{r}}^{\times})$ of $K^{\times}$ is generated by $N_{m}^{\sigma}(\xi)=\xi^{[m]_{s}}=\xi^{w}$ with $w=\gcd([m]_{s},p^{r}-1)$, and thus the order of this group is
$\frac{p^{r}-1}{w}.$
The next examples shows that using our refined notion of equivalence and isometry will give a tighter classification of constacyclic codes.
Example 5.11.
[NevPum2025, Example 5.8] Let $\gcd([m]_{s},p^{r}-1)=p^{r}-1$. Then $N_{m}^{\sigma}(\mathbb{F}_{p^{r}}^{\times})=\{1\}$ and so no two classes of skew $(\sigma,a)$-constacyclic and skew $(\sigma,b)$-constacyclic codes of length $m$ belonging to distinct $a,b\in\mathbb{F}_{p^{r}}^{\times}$ will be Chen equivalent. Thus, there are $|\mathbb{F}_{p^{r}}^{\times}|=p^{r}-1$ distinct classes of skew constacyclic codes under Chen equivalence.
Now there are at least $n$ choices for $\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})$ (corresponding to $\sigma^{\ell}$ for each $1\leq\ell<n$). This implies that the set $\{a^{p^{v}}\mid 0\leq v<r\}\subset\mathbb{F}_{q^{r}}$ is an equivalence class with $r$ elements. That is, all of the corresponding $(\sigma,a^{p^{v}})$-constacyclic codes of length $m$ are equivalent and so the set ${\bf C}_{t^{m}-a}]=\{{\bf C}_{t^{m}-a^{p^{v}}}\,|\,0\leq v<r\}$ has $r$ elements, while ${\bf C}_{t^{m}-a}]_{Chen}=\{{\bf C}_{t^{m}-a}\}$.
The following example which demonstrates how to apply the connection between skew polycyclic and skew constacyclic codes from Proposition 5.7 to a plethora of different scenarios and again clearly demonstrates that using our refined notion of equivalence gives a tighter classification of skew polycyclic codes.
Example 5.12.
Let $\gcd([m]_{s},p^{r}-1)=p^{r}-1$.
We construct the equivalence classes of some given $h(t)$ as in Theorem 5.6. We start with $h(t)=t^{m}-\sum_{i=0}^{m-1}b_{i}t^{i}$. For all $\alpha\in K^{\times}$, define
$h_{id,\alpha}(t)=t^{m}-\sum_{i=0}^{m-1}N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}t^{i}.$
Then there exists a Chen equivalence $G_{id,\alpha}$ between $R/Rh_{id,\alpha}$ and $R/Rh$ and
${\bf C}_{h}]_{Chen}=\{{\bf C}_{h_{id,\alpha}}\,|\,\alpha\in\mathbb{F}_{p^{r}}^{\times}\}.$
All the skew polynomials $h_{id,\alpha}(t)$ have constant coefficient $b_{0}$.
By Proposition 5.7, employing Example 5.11, no two classes of skew $(f,\sigma)$-polycyclic and skew $(h,\sigma)$-polycyclic codes of length $m$ belonging to $f$ and $h$ with distinct constant coefficients $a_{0},b_{0}\in K^{\times}$ will be Chen-equivalent. Thus, there are at most $|K^{\times}|=p^{r}-1$ distinct classes of skew polycyclic codes of length $m$ under Chen equivalence.
Now take $\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})$ and set
$h_{\tau,\alpha}(t)=t^{m}-\sum_{i=0}^{m-1}N_{m-i}^{\sigma}(\sigma^{i}(\tau(\alpha)))\tau(b_{i})t^{i},$
then
${\bf C}_{h}]=\{{\bf C}_{h_{\tau,\alpha}}\,|\,\alpha\in\mathbb{F}_{p^{r}}^{\times},\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})\}.$
Note that the $h_{\tau,\alpha}(t)$ all have $\tau(b_{0})$ as constant coefficients, for some $\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})$, since $N_{m}^{\sigma}(K^{\times})=\{1\}$, and that there are at least $n$ choices for $\tau$.
This argument works more generally, too.
Example 5.13.
For $i\in\{0,\ldots,m-1\}$ set $m_{i}=m-i$. Suppose there is $i\in\{0,\ldots,m-1\}$ such that $\gcd([m_{i}]_{s},p^{r}-1)=p^{r}-1$. By Proposition 5.7, no two classes of skew $(f,\sigma)$-polycyclic and skew $(h,\sigma)$-polycyclic codes of length $m$ where $f$ and $h$ have distinct $a_{i_{0}},b_{i_{0}}\in\mathbb{F}_{p^{r}}^{\times}$ will be Chen-equivalent. Thus, there are again at most $|K^{\times}|=p^{r}-1$ distinct classes of skew polycyclic codes under $m$-Chen equivalence; one for each choice of $a_{i_{0}}\in\mathbb{F}_{p^{r}}^{\times}$. Note that every ${\bf C}_{h_{id,\alpha}}\in[{\bf C}_{h}]_{Chen}$ has $b_{i_{0}}$ as $i_{0}$-th coefficient. Every ${\bf C}_{h_{\tau,\alpha}}\in[{\bf C}_{h}]$ has $\tau(b_{i_{0}})$ as $i_{0}$th coefficient.
We will use the following result for skew constacyclic codes obtained in [NevPu2025, Theorem 5.6].
Theorem 5.14.
The number of distinct Chen isometry classes of families of skew $(\sigma,a)$-constacyclic codes arising from nonassociative algebras $\mathbb{F}_{p^{r}}[t,\sigma]/\mathbb{F}_{p^{r}}[t,\sigma](t^{m}-a)$ is $N$ where
$N=\begin{cases}\gcd([m]_{s},p^{r}-1)&\text{if $n\nmid m$; and}\\ \left(1-\frac{1}{[n]_{s}}\right)\gcd([m]_{s},p^{r}-1)&\text{if $n|m$.}\end{cases}$
There are additionally $\gcd([m]_{s},p^{r}-1)/[n]_{s}$ different Chen equivalence classes (and thus at most this many Chen isometry classes) of families of skew $(\sigma,a)$-constacyclic codes arising from associative algebras $\mathbb{F}_{p^{r}}[t,\sigma]/\mathbb{F}_{p^{r}}[t,\sigma](t^{m}-a)$, that is, for which $n\mid m$ and $a\in K_{0}$.
Note that when $n|m$, we have $N_{m}^{\sigma}(\mathbb{F}_{p^{r}})\subset\mathbb{F}_{p^{s}}$.
As a consequence we immediately obtain bounds for skew polycyclic codes.
Theorem 5.15.
(i) Let $n\nmid m$. Then the number of different Chen equivalence classes of skew $(f,\sigma)$-polycyclic codes of length $m$ is at most
$N=\gcd([m]_{s},p^{r}-1).$
(ii) Let $n|m$. Then the number of different Chen equivalence classes of skew $(f,\sigma)$-polycyclic codes of length $m$ is at most
$N=\left(1-\frac{1}{[n]_{s}}\right)\gcd([m]_{s},p^{r}-1)+\gcd([m]_{s},p^{r}-1)/[n]_{s}.$
In this count, $\left(1-\frac{1}{[n]_{s}}\right)\gcd([m]_{s},p^{r}-1)$ equivalence classes belong to skew polynomials $f$ that have constant coefficient $a_{0}\notin\mathbb{F}_{p^{s}}$.
Proof.
$(i)$ If $n\nmid m$, then the algebras $\mathbb{F}_{p^{r}}[t;\sigma]/\mathbb{F}_{p^{r}}[t;\sigma](t^{m}-a)$ are not associative and there are $\gcd([m]_{s},p^{r}-1)$ classes of Chen equivalent $(\sigma,a)$-constacyclic codes. This immediately yields the assertion.
$(ii)$ If $n|m$, then $N_{m}^{\sigma}(\mathbb{F}_{p^{r}})\subset\mathbb{F}_{p^{s}}$, so that every coset of $N_{m}^{\sigma}(\mathbb{F}_{p^{r}})$ is either entirely contained in $\mathbb{F}_{p^{s}}$ or entirely disjoint from it (so corresponds to a nonassociative Petit algebra). We know that $\mathbb{F}_{p^{r}}[t;\sigma]/\mathbb{F}_{p^{r}}[t;\sigma](t^{m}-a)$ is a proper nonassociative algebra if and only if $a\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p^{s}}$. Turning to the proof of [NevPu2025, Theorem 5.6], the total number of cosets contained in $\mathbb{F}_{p^{s}}$, and thus corresponding to associative algebras where $a\in\mathbb{F}_{p^{s}}$, is equal to $\frac{p^{s}-1}{(p^{r}-1)/w}=\frac{w}{[n]_{s}}$. Subtracting these $\frac{w}{[n]_{s}}$ cosets from the total number of cosets gives the number of Chen equivalence classes for nonassociative algebra that is $N=w-w/[n]_{s}$. This yields the assertions. ∎
5.3. When two classes cannot be equivalent
If we want to know when skew polycyclic codes cannot be equivalent, there are some easy conditions we can check first. We are looking at classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes of length $m$ over $S$.
Corollary 5.16.
If any of the following conditions hold, then ${\bf C}_{f}$ and ${\bf C}_{h}$ are not equivalent.
$(i)$ There exists some $i\in\{0,\ldots,m-1\}$ such that $a_{i}=0$ but $b_{i}\not=0$ or vice versa (in other words, $(a_{0},\dots,a_{m-1})$ and $(b_{0},\dots,b_{m-1})$do not have the same Hamming weight).
$(ii)$ There exists some $i\in\{0,\ldots,m-1\}$ such that $a_{i}$ is invertible but $b_{i}$ is not, or vice versa.
This generalizes [OaHA25, Lemma 4.19] which was stated for finite fields and $\tau=id$. gain $\sigma=id$ is allowed, in which case we are dealing with polycyclic codes.
Proof.
The classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes are equivalent if and only if there exists $\alpha\in S^{\times}$ such that
$\tau(a_{i})=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$
This means:
$(i)$ $a_{i}=0$ if and only if $b_{i}=0$ for all $i\in\{0,\ldots,m-1\}$.
$(ii)$ $a_{i}$ is invertible if and only if $b_{i}$ is invertible for all $i\in\{0,\ldots,m-1\}$. ($(i)$ and $(ii)$ also follow from Proposition 4.12.) ∎
Corollary 5.17.
If any of the following conditions hold, then ${\bf C}_{f}$ and ${\bf C}_{h}$ are not equivalent.
$(i)$ $\tau(a_{m-1})\not=\sigma^{m-1}(\alpha)b_{m-1}$ for all $\alpha\in S^{\times}$ and all $\tau\in{\rm Aut}(S)$ commuting with $\sigma$.
$(ii)$ There exists some $i\in\{0,\ldots,m-1\}$ such that $a_{i}$ is invertible and
$N_{S/S_{0}}(\tau(a_{i})b_{i}^{-1})\not\in(N_{S/S_{0}}(S^{\times}))^{m-i}$
for all $\tau\in{\rm Aut}(S)$ that commute with $\sigma$.
Proof.
The classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes are equivalent if and only if there exists $\alpha\in S^{\times}$ such that
$\tau(a_{i})=N_{m-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}.$
This means:
$(i)$ $\tau(a_{m-1})=\sigma^{m-1}(\alpha)b_{m-1}$.
$(ii)$ If we apply $N_{S/S_{0}}$ to both sides of the equation we obtain
$N_{S/S_{0}}(\tau(a_{i}))=N_{S/S_{0}}(\alpha)^{m-i}N_{S/S_{0}}(b_{i})$
which is equivalent to
$N_{S/S_{0}}(\tau(a_{i})b_{i}^{-1})\in(N_{S/S_{0}}(S^{\times}))^{m-i}\subset(S_{0}^{\times})^{m-i}.$
5.4. Conclusions and further work
We initiated the study of skew polycyclic codes up to equivalence, identified skew polycyclic codes which are the same in terms of protecting the underlying information, and obtained some catalogues of codes up to isometries and equivalences with tools for de-duplicating them, identifying gaps on which further research may be focused. Our equivalence relations yield tighter classifications of isometric and equivalent codes that previously used ones.
Our next goal is to now look at special $S$, like finite fields or rings, and explicitly compute the equivalence classes. The isometry classes for skew polycyclic codes seem to be hard to find, however, the special case of skew constacyclic codes seems tractable.
It was pointed out already in [BouUl2014] that even when $S$ is a finite field, codes constructed with some $f\in S[t;\sigma,\delta_{\beta}]$ where $\delta_{\beta}(a)=\beta(\sigma(a)-a)$ is nonzero, can have a better minimum Hamming distance than those constructed when $\delta=0$. For finite fields, it is well known that $\mathbb{F}_{q}[t;\sigma,\delta_{\beta}]\cong\mathbb{F}_{q}[t;\sigma]$. However, this isomorphism of noncommutative rings changes the Hamming weight of corresponding skew $(\sigma,\delta)$-polycyclic codes, and hence their performance. Most of the current literature, however, only considers the case that $S$ is a finite field or Galois ring, and $\delta=0$, e.g. [Bha2012].
We plan to investigate the Hamming weight preserving isomorphisms, in particular $G_{\tau,\alpha}$, also when $\delta\not=0$ in a future paper. We also plan to look at the case where $S$ is not commutative.
Acknowledgments
This paper was written during the second author’s stay as a Simons Professor in Residence at the University of Ottawa. She gratefully acknowledges the support of CRM and the Simons Foundation.